### v1

# XTransfer: Cross-Modality Model Transfer for Human Sensing with Few Data at the Edge

Abstract
Deep learning for human sensing on edge systems offers significant opportunities for smart applications. However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems. Current methods that rely on transferring pre-trained models often encounter issues such as modality shift and high resource demands, resulting in substantial accuracy loss, resource overhead, and poor adaptability across different sensing applications. In this paper, we propose XTransfer, a first-of-its-kind method for resource-efficient, modality-agnostic model transfer. XTransfer freely leverages single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely repairs modality shift in pre-trained model layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models. We benchmark various baselines across diverse human sensing datasets spanning different modalities. Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance on human sensing tasks while significantly reducing the costs of sensor data collection, model training, and edge deployment.
1 Introduction
Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (humansensing2023, ; Ystgaard2023, ). With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed2018, ). Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Xiaoqiang2019, ; zhang2020, ; Yanjiao2020, ). However, existing DL solutions are often data- and resource-intensive for training and deployment (ondevicedl, ), posing significant challenges for data collection and edge deployment in human sensing (surveyiot, ; iotdatachallenge, ).
Unlike other data modalities (e.g., text or image), collecting human sensing data for training can be costly and even impractical. This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane2015, ; Jeyakumar19, ; Teh2020, ). Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Kroger2019, ; zhangipsn, ). Manual annotation of large volumes of sensor data introduces additional complexity and incurs high costs (Song2023fewshot, ; chen2020deep, ). Furthermore, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), thereby adding complexity to model training, adaptation, and edge deployment (humansurvey, ; humansensing2023, ).
Prior research has explored Few-shot Learning (FSL) (metasense, ; Wang2024, ), transfer learning (tlhar, ), and their combination (i.e., cross-domain FSL) (crossHAR, ; FewSense, ) to enable DL-based human sensing with few data. However, they either still require large-scale labeled datasets from the same singular modality to costly train source models from scratch (Song2023fewshot, ), or face challenges such as modality shift and limited adaptability when leveraging pre-trained models (Jaehoon2022, ; Ran2022, ; Sommelier2022, ), particularly in cross-modality settings.
These limitations highlight the need for a scalable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data. Recent advances in multi-modality learning (mmfi, ; chen2025xfi, ; chen2024probabilistic, ; madaan2024jointly, ; yang2024facilitating, ; baldenweg2024multimodal, ) with methods such as knowledge distillation (e.g., image-to-sensor transfer (zhao2018through, ; Xingzhe2022, ; Sevgi2020, ; Yu2023, )) and multi-modal model (e.g., CLIP-based models (girdhar2023imagebind, ; wu2023nextgpt, )), demonstrate the growing feasibility of cross-modality model transfer, driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (PyTorchhub, )), which opens new opportunities for real-world human sensing. However, due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Sommelier2022, ; Zhang2023, ; Jaehoon2022, ). Our experiments reveal that state-of-the-art baselines (Tab. 1), though effective on vision and NLP tasks, fail to reach oracle performance (metasense, ) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Fig. 2(a). These highlight the underlying challenge of latent feature misalignment across modalities. Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Sicheng2020, ; Xiangyumfs, ; NEURIPS20196048ff4e, ). While methods such as neural architecture search (NAS) (searchinsights, ; AdaptiveNet2023, ; LegoDNN2021, ), pruning (Anwar2017, ; frankle2018the, ; pers2020composition, ), or quantization (Polino2018ModelCV, ), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data. As shown in Fig. 2(c), our experiments using structural pruning (frankle2018the, ; NEURIPS2022, ) demonstrate significant model accuracy loss (50.5% on average). These challenges point to a critical gap in developing adaptable and scalable methods.
To overcome the challenges, we propose XTransfer, a first-of-its-kind cross(X)-modality model Transfer method that freely leverages public source models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks, significantly reducing the need for extensive sensor data collection and training from scratch. It also offers resource-efficient designs to optimize edge deployment, paving the way for scalable real-world applications. XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained source models under cross-modality FSL settings (see App. A), resulting in compact and adaptable models. To mitigate modality shift and prevent overfitting, it enables model repairing achieved by a novel Splice-Repair-Removal (SRR) pipeline (c.f. Sec. 4). This pipeline transforms a few sensor data to broadly adapt pre-trained model layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities. XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (c.f. Sec. 5) to identify and recombine layers of interest by conducting a layer-wise search. Building on the insight that not all repaired layers contribute, this mechanism discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (searchinsights, ). To further accelerate the search when dealing with multiple sources, it integrates a pre-search check strategy to efficiently avoid unnecessary repairing.
Comprehensive evaluations (c.f. Sec. 6) are conducted on 8 public datasets along with their corresponding pre-trained source models, spanning diverse modalities (e.g., image, text, audio, and sensing), 4 collected target sensing datasets in real-world settings, 2 public sensing datasets and 1 standard image dataset as targets (Tab. 2) on 3 commercial edge devices. The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (c.f. Sec. 6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while using few data with broad adaptability. XTransfer also streamlines the layer search process, reducing search time by 2.1 to 4 times. The main contributions of this work are:
  • A first-of-its-kind method, XTransfer, that enables resource-efficient, modality-agnostic model transfer for human sensing with few data at the edge, as shown in Fig. 3.
  • A novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities.
  • A novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner, creating compact models.
  • Extensive experimental evaluations, benchmarking state-of-the-art baselines based on various source and target datasets, validating competitive scalable performance of XTransfer.
Figure 1: Preliminary study. (a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet and HHAR dataset as target under a 5-shot setting (see App. A).(overfitting, ). (b) shows the average similarity and FSL difficulty across all sensing datasets (Tab. 2) to each source modality (e.g., Image, Text, Sensing) (c.f. Sec. 3). 2 distinct areas represent similarity levels (A–hard, B–normal). Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Sec. 4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability (c.f. Sec. 4). (c) shows model optimisation bottlenecks.
2 Related Work
Learning with few data for human sensing. Recent works in FSL and cross-domain FSL have proposed single- and multi-source methods. Single-source methods align latent features between source models and target data via distance-based loss functions (Zhao2021, ; Yue2021Prototypical, ), whereas multi-source methods leverage the Wisdom of the Crowd principle through source distillation (Sicheng2020, ; Xiangyumfs, ) and maximal correlation analysis (NEURIPS20196048ff4e, ). Existing attempts on human sensing tasks (metasense, ; Wang2024, ; tlhar, ; crossHAR, ; FewSense, ) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and model training with limited adaptability. Recent efforts have also explored fine-tuning large language models (e.g., GPT2 (zhou2023one, )) for time-series analysis. However, testing on benchmarks (Boosting2021, ; Jaehoon2022, ; chen2018a, ), our results reveal that they still suffer from severe overfitting under modality shift and few data. Also, none of them explicitly considers resource constraints of edge systems. In contrast, XTransfer enables model repairing to safely mitigate modality shifts using only few sensor data, enabling modality-agnostic model transfer with broad adaptability.
Resource constraints optimization. Model compression techniques such as model pruning (Anwar2017, ; frankle2018the, ; pers2020composition, ), quantization using low-precision data types (Polino2018ModelCV, ), model merging by fine-tuning shared layers (Gemel2023, ), one-size-fits-all model (cai2020once, ), and NAS techniques such as model structure optimization (AdaptiveNet2023, ; LegoDNN2021, ) or search using few data (neuralFT, ; xu2022fewshot, ), have shown success in reducing model resource overhead. However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance. XTransfer uniquely applies layer recombining to optimize model structures by efficiently searching and recombining layers of interest from pre-trained source models. It not only creates compact models for streamlined edge deployment but also further enhances model repairing.
3 Preliminary
Insights in cross-modality. We apply a standard cross-domain FSL benchmarking (Jaehoon2022, ) to gain insights. It uses the Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets, including 5 human sensing target datasets and 2 well-known image datasets (Tab. 2). It also assesses FSL difficulty by fine-tuning source models on targets. Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target. (e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities. The key insights are shown in Fig. 2(b).
Layer-wise analysis. We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer. Since effective training relies on learning discriminative features (islam2021broad, ), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (i.e., channel importance) (pmlrv162luo22c, ) of each layer to indicate layer-wise performance. We define a single layer or a dependent layer block (i.e., L-Blocks in Tab. 3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit). Fig. 2(a) shows that negative transfer (Zhang2023, ) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops. These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Fig. 2(b), indicating inconsistent latent feature interpretation during fine-tuning. To track MMC shifts, we essentially employ the Silhouette score (S-score) (shahapure2020cluster, ), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
Challenges. We aim to minimize layer-wise MMC shifts, but several technical challenges arise. First, the pre-trained feature space at each layer is highly complex due to the multidimensional nature of MMC, and this complexity increases with layer depth. Such complexity poses challenges for "repairing" latent feature distributions across modalities using few sensor data. Second, strong layer-wise dependencies, formed during source model training, become increasingly rigid in deeper layers, posing challenges for standard techniques (e.g., N-Pair loss (Sohn2016, ), Triplet loss (Schroff2015, )) that may fail to preserve these dependencies. Third, Fig. 2(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence, indicating that not all repaired layers contribute positively. This motivates our LWS control design, which selectively recombines effective layers (i.e., layers of interest) while discarding others. However, such selection must account for on-device resource constraints to optimize the layer growth. The search space also expands rapidly with the number of layers or source models, further complicating efficient search.
Figure 2: Design insights. (a) Layer-wise accuracy convergence using baselines is disrupted due to modality shift. (b) A notable MMC shift emerges and grows with increasing layer index, i.e., accuracy increases while MMC shift drops in Area A and begins to drop as MMC shift largely grows in Area B. (c) After repairing, layer S-score improves, but stagnation occurs at certain layers.
4 Model Repairing
Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs in cross-modality settings. SRR pipeline involves three key stages shown in Fig. 3. In the Splice stage, input shapes between source models and sensor data must be aligned well, as default reshaping (c.f. App. B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features across channel, height, and width, especially when layer shapes vary during layer-wise search. To enhance, we design a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structures. In the next Repair stage, each connector is fine-tuned to generate adaptive channels, preserve key sensing features, and minimize MMC shift via the generative transfer module. In the following Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage. For more design details, see Appendix B.2.
Figure 3: XTransfer overview. LWS control segments source models into layers and uses the pre-search check to decide if repairing is needed. SRR pipeline then fine-tunes connectors to repair selected layers. Finally, LWS control selects and recombines layers of interest into a compact model.
4.1 Repairing in complex feature space
Since the original latent features from source models are well-trained for high discriminability (c.f. Sec. 3), our intuition is to safely fine-tune the misinterpreted latent features of few sensor data by aligning with the original features (i.e., anchors) to minimize MMC shifts. We hence propose anchor-based cross-modality supervision alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs. To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant ones.
Reduced-dimensional feature space. We project anchor MMCs into a lower-dimensional feature space using principal component analysis (PCA) (VALPOLA2015143, ) (i.e., anchor PCA space). PCA enables a lower-dimensional subspace (i.e., adjusted by the number of PCA components, with 2 components used by default) where the projected data exhibits the highest possible variance (i.e., highly clustered). The optimized component weights (i.e., projection coefficients) highlight the most important channels that contribute to the performance of each layer. Given high accuracy achieved in each original layer, the projected anchor MMCs show high clustering performance, e.g., the centroid of the distribution of each class (i.e., Anchor-center) clustered tightly, as shown in Fig. 3(a). Once the anchor PCA space is initialized for each layer, we project sensing MMCs into its space using the same parameters.
Feature space alignment. After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Fig. 3(a). Specifically, the scale and orientation of Sensing-center are notably different from those of Anchor-center, caused by MMC shifts. To align the scale, we first define a scale function based on the mean shift in inter-class distance $\mathrm{InterD}(\cdot)$, formulated as $S={{Mean}({InterD}(Pro(fs_{ij})))}\mathbin{/}{{Mean}({InterD}(Pro(ft_{ij})))}$, where $Pro(\cdot)={PCA}({MMC}(\cdot))$ denotes the space project function, $fs$ and $ft$ represent the latent features from both source and sensing modalities, respectively. $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$. Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (nguyen2010cosine, ) by fine-tuning a multi-dimensional rotation matrix $[M_{rot}]$ (where the dimensions equal to the number of PCA components) as below:
$\footnotesize\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
where $Cos(\cdot)$ denotes the cosine similarity function, and $Cen(\cdot)$ calculates the centroid of the projected distribution. Fig. 3(b) shows the aligned results.
Anchor class pairing. Building on key finding that class pairing boosts latent feature alignment across modalities (c.f. Sec. 3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the supervision alignment. We minimize pairing shift (i.e., the sum distance across each pair) via a linear sum assignment solved using the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
4.2 Repairing layer-wise dependence
Generative transfer module. To effectively restore layer-wise dependence, we design a generative transfer module that enhances the supervision alignment through generative learning. It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Fig. 3). Freezing preserves stable anchors and prevents overfitting. For each layer, given the optimal pairing set $pairST$, the generator is fine-tuned using the discriminator’s loss to minimize MMC shifts, and uniquely transforms sensor inputs to align with anchors.
Anchor-based repair loss. Building on the generative transfer module, we propose an anchor-based repair loss function ($loss_{srr}$) to harmonize latent feature distributions across modalities. Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the Euclidean distance (ED) ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss). Since the projected anchor MMCs are highly clustered, it can provide the maximum margin ($M_{max}$) between anchor clusters. To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance. Hence, our loss function is formulated as:
$\footnotesize\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
where $K$ denotes the number of negative samples, and $N$ denotes the number of classes. The $ReLU(\cdot)$ is used to trigger the negative loss in the condition where the distance between the samples is less than $M_{max}$. Fig. 3(c) shows the effectiveness of model repairing.
5 Layer Recombining
Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) efficiently select and recombine layers of interest to form a compact model.
5.1 Search control design
Following NAS principles (searchinsights, ), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency. As layers of interest grow, resource overhead can increase exponentially, making dynamic search control essential under on-device resource constraints. To efficiently manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$. Search actions (init, continue, skip, cross) are defined to guide layer recombination within each designated search pool. Init selects the starting layer with a default fixed depth of 1. Continue and skip guide intra-model layer recombination, while cross enables inter-model layer recombination.
Search strategy under resource constraints. To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate both individual repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$. It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ denotes the connector and $L_{ij}$ represents each repaired layer. Under resource constraints, only the top-performing layer in each pool is selected. LWS control then checks whether its S-score exceeds that of the previous layer. If successful, the layer $L_{ij}$ is identified as the layer of interest and recombined, and the next pool window begins at layer index $j+1$. Otherwise, the current pool is discarded, and the search window shifts forward. Considering the constraints, we design a resource-constrained function that works in conjunction with $V_{ij}(\cdot)$, which is formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~{}L]$, where $n$ denotes the current recombined layer index, and $L$ denotes the maximum number of L-units. Each layer’s adjusted overhead, $R(n)_{ij}$, is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ calculates actual cost (e.g., FLOPs), and $Max(Res(P))$ is the peak cost in current pool $P$. The final resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$. The objective is then defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is the total allowed resource budget.
5.2 Efficient search
Pre-search check. LWS control is designed to handle large search spaces from deep or multi-source models. While fine-tuning the connectors in SRR pipeline is efficient, the large search space can considerably slow down LWS control. We hence propose the pre-search check strategy to avoid unnecessary repair steps before applying SRR. We track S-score progression across three states—after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After) (i.e., repaired). SRR pipeline consistently improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation. Based on this, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs. Experiments show the actual repair rate grows exponentially with the recombined layer index (see App. B.4). We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective function defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers. The estimation begins after init. Once a new layer is selected, LWS control updates the estimation. Due to the limited updates, the estimation may be inaccurate. To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in App. B.5).
6 Experiment
6.1 Experiment setup
Table 1: State-of-the-art baselines.
  Baselines Group Description
  TL (TransferabilityID, ) Single-source A standard transfer learning method by freezing all layers
FT (Jaehoon2022, ) Single-source A standard fine-tuning method by updating all layers
ProtoNet (Snell2017, ) Single-source A distance-based FSL method
MAML (Finn2017, ) Single-source A learning-based FSL method
DAPN (Zhao2021, ) Single-source A cross-domain FSL method using adversarial ProtoNet
GPT2 (zhou2023one, ) Single-(LLM) A large language model based method using fine-tuning
MDDA (Sicheng2020, ) Multi-source A cross-domain FSL method using knowledge distillation
MCW (NEURIPS20196048ff4e, ) Multi-source A cross-domain FSL method using maximal correlation analysis
MetaSense (metasense, ) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
Table 2: Source and target datasets specifications.
  Dataset Modality Subject Class Type Shape Privacy
  Source
miniImageNet (miniimagenet, ) Image - 100 - 3x84x84 Public
miniDomainNet (domainnet, ) Image - 126 - 3x224x224 Public
Office-31 (office31, ) Image - 31 - 3x224x224 Public
Office-Home (officehome, ) Image - 65 - 3x224x224 Public
Caltech-101 (caltech101, ) Image - 101 - 3x224x224 Public
Newsgroup (text2021, ) Text - 20 - 100x1x1000 Public
VoxCeleb (VoxCeleb2017, ) Audio - 100 Time-S 1x512x300 Public
OPPORTUNITY (Roggen2010, ) IMU 12 11 Time-S 77x1x100 Public
Target
HHAR (Stisen2015, ) IMU 9 6 Time-S 6x1x256 Public
WESAD (Schmidt2018, ) ECG 15 3 Time-S 10x1x256 Public
Gesture IMU/PPG 10 8 Time-S 7x1x200 Private
Writing Ultrasound 10 10 Time-S 1x1x256 Private
Emotion mmWave 10 7 Spectrum 3x224x224 Private
BP mmWave 10 - Doppler 1x1x499 Private
ChestX (Wang2017, ) Image - 5 - 1x1024x1024 Public
  Backbone Layers L-Blocks GFLOPs Params Kernel Shape
  ResNet18 (chen2018a, ) 18 9 3.67 11.18M 2D 3x224x224
ResNet18-1D (chen2018a, ) 18 9 0.39 3.9M 1D 100x1x1000
Multi-ResNet18 (Sicheng2020, ) 18x5 9 18.35 55.9M 2D 3x224x224
Conv4 (chen2018a, ) 4 4 0.02 0.07 1D 1x1x500
Conv5 (metasense, ) 5 5 0.19 30.42M 1D By Input
GPT2 (zhou2023one, ) 4x6 6 0.41 82.9M 1D By Input
Table 3: DL model backbone specifications.
We benchmark XTransfer against state-of-the-art baselines, grouped into single- and multi-source methods (Tab. 1). Evaluation is conducted on 8 widely-used public datasets for cross-modality model transfer (Tab. 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with standard backbones such as ResNet18 from PyTorch Hub (PyTorchhub, ) (Tab. 3). To evaluate XTransfer, we develop 4 real-world applications and build three testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, raspberry pi). Based on these, we collect 4 human sensing datasets (marked as Private in Tab. 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute. Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking. We apply standard Leave-One-Out Cross-Validation (LOOCV) (metasense, ) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default. For more details, see Appendix A.
6.2 XTransfer performance
Figure 4: Ablation study evaluating the performance of model repairing and layer recombining.
Repair performance. We first evaluate how our Repair loss reduces MMC shift to enhance the S-score of each layer with different MMC dimensionalities in the SRR pipeline. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting. The channel removal is disabled in this experiment. We test with 4 different levels of MMC dimensionality (i.e., number of channels) ranging from 64 to 512. We employ N-Pair loss (Sohn2016, ) and Triplet loss (Schroff2015, ) as baselines. We run 15 rounds for each setting by default throughout the evaluation. Fig. 4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality. It indicates that Repair loss effectively enhances layer performance.
Efficient search performance. We now evaluate LWS control performance regarding model accuracy and layer search efficiency. We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Tab. 2, with a default search depth of 3 in 3-, 5- and 10-shot settings. We also utilize MetaSense in Tab. 1 as the oracle baseline. We compare different configurations: ResNet18 (Single), Multi-ResNet18 (Multi) without efficient search enabled, and Multi-ResNet18 (Multi-Efficient) with efficient search enabled. As expected, Fig. 4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline. Additionally, Multi-Efficient improves average accuracy by 7.8% over Multi across the three settings. While Multi-Efficient takes a longer search time than Single, it successfully achieves superior accuracy, highlighting the benefits of using multiple source models in XTransfer. The results also reveal that both Multi-Efficient and Multi achieve lower standard deviation on average compared to Single, suggesting that using multiple source models offers commendable search stability. In short, the results prove that the proposed efficient search significantly accelerates the layer search process and achieves better model accuracy.
Impact of different sources. Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models. We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Tab. 2. We keep the other experimental setups unchanged. Fig. 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively. This suggests that the latent features learned from Image have better discriminability than those from other modalities. Since XTransfer prioritizes accuracy, we select Image source models by default. More ablation study results refer to Appendix C.
6.3 Comprehensive comparison
Cross-modality model transfer performance. We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets, as shown in Tab. 2. We compare XTransfer against a range of state-of-the-art baselines shown in Tab. 1. We also mark our methods as Our-Single and Our-Multi, using pre-trained ResNet18 on miniImageNet and Multi-ResNet18 with 5 corresponding image datasets as sources. In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (zhou2023one, ). All methods on target datasets are evaluated under the same cross-modality FSL settings (see App. A), with MetaSense (metasense, ) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Tab. 3). For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels. As shown in Tab.4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings. Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively. Interestingly, in 3-shot setting, the averaged model accuracy of our methods is slightly lower than that of the oracle on the HHAR, Gesture, and Emotion datasets. This suggests that XTransfer’s Repair loss may lead to less stable improvements for latent feature alignment when using extremely low sample sizes, but its performance remains strong and stable in 5- and 10-shot settings. In summary, the comprehensive results show that XTransfer achieves state-of-the-art model accuracy, substantially reducing the cost of data collection and model resource overhead for scalable deployment of human sensing applications.
Training resource usage. We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) on the cloud. We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Finn2017, ) and DAPN (Zhao2021, )), excluding methods that rely solely on fine-tuning or distance-based techniques. The experimental setup remains the same as above. As shown in Tab. 6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines. Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines. This is due to our layer-wise method, which progressively repairs and recombines only a subset of layers rather than training the entire model at once. It also highlights the effectiveness of our compact connector design. Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency. Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer with low computational overhead.
HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
MetaSense (metasense, ) 63.2 69.0 75.2 55.1 64.0 68.3 66.3 73.4 79.4 72.1 83.3 89.6 55.5 56.3 65.5 24.8 28.1 31.7
ProtoNet (Snell2017, ) 50.6 45.7 49.8 43.9 49.3 52.0 50.8 55.8 59.2 69.5 78.7 73.4 51.3 49.2 54.8 21.7 23.4 24.5
DAPN (Zhao2021, ) 48.7 51.2 54.7 56.9 61.2 63.1 45.8 49.4 50.5 75.1 78.6 79.6 48.0 50.0 58.7 24.0 24.4 25.5
MAML (Finn2017, ) 36.6 42.3 42.4 39.6 57.9 65.6 37.2 41.3 39.4 31.6 39.7 35.2 22.3 26.9 31.1 22.5 20.4 21.9
GPT2 (zhou2023one, ) 41.0 34.0 45.0 48.0 35.0 36.0 49.0 54.0 63.0 52.0 71.0 75.0 - - - - - -
Our-Single 59.0 71.8 80.7 77.9 78.4 81.8 45.2 69.6 77.8 72.8 87.0 91.3 43.6 55.6 61.4 27.7 28.6 30.4
MDDA (Sicheng2020, ) 13.2 17.7 18.8 40.0 39.3 41.9 11.2 13.9 15.2 12.5 10.9 9.0 14.4 15.3 16.0 20.6 19.6 19.9
MCW (NEURIPS20196048ff4e, ) 27.6 25.4 26.9 30.7 32.7 32.4 22.3 23.4 25.6 32.3 31.6 30.2 25.3 27.2 26.5 22.4 21.5 22.8
Our-Multi 59.3 74.3 78.4 76.4 77.8 78.7 48.7 73.1 78.6 74.1 86.1 90.4 42.3 55.1 58.6 29.0 30.0 30.2
Table 4: Comparison in model accuracy. Top-2 results are highlighted in bold.
HHAR WESAD Gesture Writing Emotion ChestX
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
ProtoNet (Snell2017, ) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
DAPN (Zhao2021, ) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
MAML (Finn2017, ) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
GPT2 (zhou2023one, ) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
Our-Single 1.42 1.57 2.09 2.76 2.61 2.51 1.55 1.90 1.76 2.11 1.88 1.84 0.82 1.07 1.26 0.83 1.04 1.24
MDDA (Sicheng2020, ) 0.03 0.04 0.04 0.08 0.08 0.08 0.02 0.03 0.03 0.03 0.02 0.02 0.03 0.03 0.03 0.04 0.04 0.04
MCW (NEURIPS20196048ff4e, ) 0.06 0.05 0.05 0.06 0.07 0.06 0.04 0.05 0.05 0.06 0.06 0.06 0.05 0.05 0.05 0.04 0.04 0.05
Our-Multi 1.71 1.47 1.64 3.00 2.70 2.75 1.20 1.25 1.63 1.93 1.96 1.53 0.96 1.23 1.73 0.68 1.24 0.83
Table 5: Comparison in ATR ratio. Top-2 results are highlighted in bold
HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
DAPN (Zhao2021, ) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
MAML (Finn2017, ) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
Our-Single 18.27 18.90 19.48 15.95 16.46 16.98 20.67 21.71 22.51 19.92 20.67 21.08 18.24 19.20 19.96 16.97 17.58 18.80
Our-Multi 18.31 20.77 22.06 18.00 19.42 23.10 20.49 22.49 24.49 18.68 19.25 21.69 16.45 18.18 19.85 15.33 17.69 19.62
Table 6: Comparison in GPU utilization of each method during training. Top-2 results are highlighted in bold
7 Conclusion
This paper presents XTransfer, a novel method that enables resource-efficient, modality-agnostic model transfer for human sensing applications at the edge. XTransfer harnesses the power of pre-trained models as free sources and requires only few sensor data to create compact models. It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a practical and scalable method for advancing human sensing.
Limitations and future work. Since XTransfer reuses layers from one or more pre-trained source models, its performance depends on their quality (c.f. Sec. 4). While performance varies with different source models, XTransfer’s modality-agnostic design ensures broad adaptability, enabling flexible use of diverse source models without concern for modality shift or resource constraints. For fair benchmarking, we primarily use the pre-trained ResNet18 from PyTorch Hub (PyTorchhub, ) (c.f. Sec. 6.1), but XTransfer’s layer-wise repairing and recombining support flexible backbone choices. While current version requires source models with the same backbone, future versions will support mixed backbones to further enhance flexibility.
Appendix
The appendix is organized as follows:
  • $\bullet$
    Section A illustrates more details on our experiment setup, including implementation details, developed real-world sensing applications, experimental setups, dataset statistics, evaluation metrics, and computation resources we used for our experiments.
  • $\bullet$
    Section B introduces additional technique details of different components in XTransfer.
  • $\bullet$
    Section C gives more experimental results about ablation study and benchmarking.
Appendix A Experiment setup details
A.1 Settings for learning with few data
A typical FSL method consists of two stages–meta-training and meta-testing [56]. In meta-training stage, FSL requires a large-scale source dataset from the same singular modality (i.e., typically specific to the application) to train source models from scratch. The source dataset is divided into a support set and a query set  [^4] The support set and query set are assumed to be drawn from the same data distribution [82].. Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format [15]. The meta-testing stage follows the same structure but uses unseen target data. Specifically, the target support set is used to adapt the source models, which are then evaluated on the target query set. Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing [18]. LOOCV works by leaving one user out as the target for meta-testing while using the remaining users to train the source models during meta-training in each round. The MetaSense [18] is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage. Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same singular modality [23]. In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but with pre-trained source models selected across different modalities without constraints.
A.2 Setups in preliminary study
As shown in Fig. 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ [51]. In Fig. 2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub [38]) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Tab. 2. We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Tab. 2. We use a default 5-shot cross-modality FSL setting with LOOCV enabled.
A.3 Setups in layer-wise analysis
In Section 3, we conduct a layer-wise analysis to gain insights into cross-modality model transfer with few sensor data. To set up, we continue to use the standard ResNet18 [38] and segment it as 9 L-Blocks, as shown in Tab. 3. We also employ the default reshaping (c.f. App. B.1) on HHAR as target with 5-shot setting, and use the original source model input (miniImageNet) as reference, marked as Original. We examine the single-source baselines including Transfer Learning (TL) [68], Fine-tuning (FT) [23], MAML [70], DAPN [52], and MetaSense [18] as the oracle baseline, shown in Tab. 1. To precisely track the MMC shift changes at each layer during fine-tuning,
A.4 Layer-wise metric
To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) [63] as the layer-wise metric to evaluate how well the features are being fine-tuned. The S-score is in the range from -1 to 1 and being better if closing to 1. To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL [68] and Original methods. Fig. 6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner. TL’s S-score also performs equivalently to its accuracy (i.e., S-score drop along with accuracy drop) in detecting the negative transfer across different layers.
A.5 Implementation
We implement XTransfer using PyTorch version 1.4.0. We also benchmark XTransfer with a full development cycle of cross-modality FSL tasks, including pre- and post-processing, training, and inference on commercial edge devices. Once pre-trained source models and few sensor data are processed in cloud via XTransfer, compact models can be seamlessly converted to various DL platforms (e.g., PyTorch Mobile, TFLite) for broader edge deployments. To enable reproducibility of the results, we fix all random seeds in the code, set the evaluation mode consistently when evaluating output models, and enable CUDA convolution determinism. To measure model resource overhead (e.g., FLOPs and parameters), we use the standard PyTorch OpCounter. For regression tasks, class-based transformation [83] enables compatibility.
A.6 Developed sensing applications
To evaluate XTransfer on real-world human sensing targets, we essentially develop 4 human sensing applications, as shown in Fig. 5. These include: (1) emotion recognition via facial expressions (i.e., Emotion) and (2) non-contact blood pressure monitoring (i.e., BP) based on an embedded mmWave radar setup, (3) handwriting tracking (i.e., Writing) using ultrasound based on a smartphone, and (4) gesture recognition (i.e., Gesture) using IMU and PPG sensors based on a smartwatch. The data collection is approved with ethics approval and involves 40 subjects. The corresponding target sensing datasets are shown in Tab. 2 with detailed specification.
A.7 Hardware setup and computation environment
We evaluate XTransfer on real-world testbeds using 3 commercial edge devices, ranging from smartwatch to Raspberry Pi (Tab. 7). For example, Fig. 5(a) presents an embedded mmWave radar setup which consists of a TI IWR1843Boost and a Raspberry Pi 4 using ROS Melodic on Ubuntu 18.04.4. In addition, we implement both smartwatch and smartphone testbeds using PyTorch Android version 1.10.0. We also employ servers equipped with NVIDIA GeForce RTX 3090 GPU to build our cloud environment.
A.8 Training details
In the Repair stage, we fine-tune the generative transfer module (c.f. Sec. 4) for 100 episodes using SGD with a learning rate of 1e-2 and momentum 0.95. We apply a step-based learning rate scheduler with a step size of 20 and a decay factor of 0.5. We also use standard early stopping with a patience of 20. We retain the training parameters of the baselines in Tab. 1 as specified in their original implementations.
Figure 5: (a) Embedded mmWave radar testbed setup; (b)-(e) Built human sensing applications across different real-world settings.
  Device ROM RAM CPU Battery OS
  Galaxy Watch 4 16GB 1.5GB Exynos W920 247mAh Wear OS 3.5
Raspberry Pi 4 64GB 8GB Broadcom BCM2711 - Ubuntu 18.04.4
Redmi K20 Pro 128GB 8GB Snapdragon 855 4000mAh Android 11
Table 7: Edge devices specification.
Appendix B Technical details
Figure 6: Design insights. (a) shows layer-wise metric correlation. (b)(c) present efficient search insights into LWS control using multiple source models.
B.1 Default reshaping
To align with source model input shape, we develop a default reshaping to transform sensor data shape. It uses bilinear interpolation [84] (i.e., Resizer) to resize the height and width, and a fixed convolutional layer (i.e., Fixed header) with a $1\times 1$ kernel to match the input channels [85], placed between source models and sensor data.
B.2 PCA-based channel removal
In the Removal stage, channel removal is designed to further enhance layer performance while optimizing resource usage. This is achieved by removing unnecessary channels at each layer. Existing pruning techniques [46, 47] typically rely on channel importance metrics, such as the L2 norm, calculated from MMC. However, due to MMC shift, these metrics may inaccurately assess channel importance, resulting in degraded layer performance. To address this, we propose using PCA-based channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for channel removal. Our objective is to remove unnecessary channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
B.3 Setups for efficient search insights
To gain design insights of efficient search for LWS control (c.f. Sec. 5), we obverse a complete process where all layers are repaired using the HHAR dataset and a Multi-ResNet18 model (i.e., 5 source models with 45 L-unit), as shown in Tab. 3. Fig. 6(b) shows the changes in the S-score across three states: after pairing anchors (i.e., Anchor), before and after applying the SRR pipeline (i.e., Before and After, respectively), in a layer-wise manner. The results show that the SRR pipeline successfully increases the S-score from Before state (i.e., lower bound) to align with the Anchor state (i.e., the objective). This indicates a strong correlation between the After state’s S-score and both Before and Anchor values. These insights lead to the observation for designing the repair rate growth model.
B.4 Observation for designing repair rate growth model
Our key idea is to estimate the repaired S-score (i.e., repair value) of each layer before initiating SRR processing, allowing us to decide whether to bypass unnecessary repairs. To empirically examine this, we first define a repair rate function to capture this correlation as $rate_{n}=(After-Before)/Anchor=(V_{ij}^{n}-Score(Pro(ft_{ij}^{n}),Y_{t}))/Score(Pro(fs_{ij}^{n}),Y_{s})$, where $Y_{s}$ denotes the source labels, and we normalize $rate_{n}\in[0,1]$. We then repeat the process multiple times to observe the changes in $rate_{n}$. Fig. 6(c) shows that the repair rate demonstrates exponential growth. Thus, we design the repair rate growth model based on an exponential growth model (c.f. Sec. 5.2).
B.5 Dynamic search range mechanism
Based on the estimated repair rate, LWS control can initially use a small, fixed search range to efficiently filter out most layer candidates before initiating SRR processing. However, because the updates to $rate_{n}^{est}$ are limited, the estimation may be inaccurate, potentially affecting selection decisions (e.g., missing the optimal layer candidate). To mitigate this risk, we develop a dynamic search range mechanism that adjusts the search range to balance accuracy and efficiency. This mechanism is formulated as $range_{n}=(1+\hat{u}^{\pm}rate_{n}^{est})*range_{n-1},\,range_{n}\in[0.2,1]$, where the range starts at 0.5 and adjusts based on the previous range. Here, $\hat{u}$ denotes a signed unit vector. If $rate_{n-1}^{est}$ is less than $rate_{n-1}$, it indicates that the estimation might be inaccurate and $\hat{u}$ will be positive to increase the range in searching for the optimal layer. Otherwise, the range will be reduced to improve search efficiency.
B.6 Post fine-tuning
After LWS control is completed, we add either a linear classifier or a regression layer (e.g., a fully connected layer) at the end of the model to finalize its backbone. To strengthen layer-wise dependence, we perform a quick post fine-tuning process, updating only the connectors and fully connected layers.
Appendix C Additional experiment results
Figure 7: Ablation study evaluating the performance of components, search parameters, and applications.
C.1 Removal performance
We evaluate how channel removal enhances the average S-score of selected layers. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting, and enable channel removal with a default search depth of 3. We employ L2 norm-based (i.e., L2-based) model pruning [47, 50] as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning). Fig. 7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers. L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the channel importance may be inaccurately estimated, resulting in degraded performance. These results show that the proposed channel removal further improves layer performance by accurately removing unnecessary channels at each layer.
C.2 Impact of search depth
In this experiment, we investigate how different search depths affect the accuracy and resource overhead (e.g., FLOPs) of output models. Varying the search depth alters the search window sizes with the number of layer candidates, which can lead to different output model backbones and accuracy. To quantify these differences, we employ HHAR as target, ResNet18 on miniImageNet (Single) and Multi-ResNet18 (Multi) on 5 image datasets as sources (shown in Tab. 2) in 3-, 5- and 10-shot settings. We also utilize MetaSense in Tab. 1 as the oracle baseline. Both Single and Multi models have up to 9 L-unit, hence we set three different search depths as D1, D3, D9. Fig. 7(b) shows that the output models of both Single-D1 and Multi-D1 result in high FLOPs, while that of both Single-D9 and Multi-D9 achieves low FLOPs. As search depth decreases, the search process becomes longer, and it triggers LWS control to make more decisions on recombining layers, which results in a suboptimal model backbone with high FLOPs. Although both Single-D3 and Multi-D3 do not produce the lowest FLOPs on average, they outperform others and achieve the highest average model accuracy of 70.52% and 70.69%, respectively, across the three settings. Therefore, we set D3 as the default search depth, as it provides the best balance between resource overhead and accuracy of the output models.
C.3 Regression task performance
We also evaluate the effectiveness of XTransfer in a regression task using the BP dataset, which includes both systolic (SBP) and diastolic (DBP) sensor data shown in Tab. 2 collected by the blood pressure monitoring application. We employ the pre-trained ResNet18 on miniImageNet as the source model and use the 5-shot setting. To integrate this into the SRR pipeline, we transform the regression task by treating individual regression labels as distinct classes for anchor class pairing. For benchmarking, we use mean error (ME) and standard deviation (STD), following the Advancement of Medical Instrumentation (AAMI) standard, which accepts an ME within 5 mmHg and STD within 8 mmHg [86]. We employ a random forest (RF) based method [87] as the oracle baseline. Besides, we train a Conv4 backbone in Tab. 3 (Conv) using LOOCV, and the fine-tuned Conv (ConvFT) as the baselines. Fig. 7(c) shows that both XTransfer and the RF methods surpass the AAMI standard. In particular, XTransfer outperforms the baselines and RF in estimating both SBP (ME: -0.64mmHg, STD: 7.2mmHg) and DBP (ME: 0.07mmHg, STD: 7.47mmHg) tasks. These results demonstrate that XTransfer is highly efficient in handling regression tasks with few sensor data.
C.4 On-device performance
We now evaluate the on-device performance of deploying XTransfer’s output models on 3 commercial edge devices (Tab. 7), focusing on resource overhead metrics such as FLOPs, parameters, latency per call, and peak memory footprint. For each target dataset, we test both types of output models created by our single-source (Single) and multi-source (Multi) methods. We also use ResNet18 and Multi-ResNet18 (Tab. 3) as baselines. As shown in Tab. 8, all output models achieve significant reductions in model resource overhead, e.g., FLOPs is reduced by 2.4 to 5.7$\times$ for Single models and by 10.1 to 16.5$\times$ for Multi models, compared to the baselines. These recombined models also demonstrate substantial improvement in on-device performance. Specifically, in the Single method, latency is reduced by 1.4 to 4.1 $\times$ and peak memory footprint decreases by 1.8 to 3.9$\times$. In the Multi method, the reductions are even more pronounced, with latency dropping by 7.4 to 21$\times$ and memory footprint by 2.2 to 8.4$\times$, compared to the baselines. These results prove that XTransfer successfully achieves streamlined edge deployment.
Watch Pi Phone
Dataset Type FLOPs Params Latency (ms) Mem (MB) Latency (ms) Mem (MB) Latency (ms) Mem (MB)
HHAR Single 1.51G 0.75M 625.5 125.5 274.18 246.62 26.86 179.5
Multi 1.46G 2.16M 634.5 140 249.26 247.83 28.34 209
WESAD Single 1.19G 0.74M 604.5 120 240.05 244.64 25.77 168
Multi 0.89G 0.34M 460.5 122 209.52 235.55 21.57 181
Gesture Single 1.21G 2.25M 706.5 135 259.25 251.26 31.5 192
Multi 1.45G 2.09M 692 137 268.79 253.7 30.96 190
Writing Single 1.32G 0.87M 467.5 126 261.71 242.33 29.79 183.5
Multi 1.28G 0.73M 513 124.5 285.2 247.01 34.67 185
Emotion Single 1.51G 0.49M 767.5 130.5 390.36 271.02 41.1 171.5
Multi 1.43G 2.14M 781 142 285.16 251.48 37.9 189.5
ChestX Single 0.64G 0.72M 690.5 126 210.45 237.45 29.57 165
Multi 1.06G 0.7M 670 130.5 261.14 236.96 30.14 183.5
ResNet18 - 3.67G 11.18M 1060 189 703.52 303.75 89.8 311
Multi-ResNet18 - 14.68G 55.9M 5816 356 4242.27 482.26 454.44 436
Table 8: On-device Performance.