### v2

# XTransfer: Modality-Agnostic Few-Shot Model Transfer for Human Sensing at the Edge

Abstract
Deep learning for human sensing on edge systems presents significant potential for smart applications. However, its training and development are hindered by the limited availability of sensor data and resource constraints of edge systems. While transferring pre-trained models to different sensing applications is promising, existing methods often require extensive sensor data and computational resources, resulting in high costs and poor adaptability in practice. In this paper, we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer with resource-efficient design. XTransfer flexibly uses single or multiple pre-trained models and transfers knowledge across different modalities by (i) model repairing that safely mitigates modality shift by adapting pre-trained layers with only few sensor data, and (ii) layer recombining that efficiently searches and recombines layers of interest from source models in a layer-wise manner to create compact models. We benchmark various baselines across diverse human sensing datasets spanning different modalities. Comprehensive results demonstrate that XTransfer achieves state-of-the-art performance while significantly reducing the costs of sensor data collection, model training, and edge deployment.
1 Introduction
Human sensing refers to the process of capturing and interpreting data related to human activities, behaviors, and physiological states using various sensors (Zhang et al., 2023a; Ystgaard et al., 2023). With the proliferation of edge devices equipped with sensors, human sensing plays a vital role in edge systems to understand contexts and enable smart applications, ranging from activity recognition to emotion or vital sign detection (Ahamed & Farid, 2018). Deep learning (DL) offers robust performance in interpreting sensor data, and its deployment on edge devices improves privacy and reduces bandwidth (Zhang et al., 2020; Chen et al., 2020). However, existing DL solutions are often data- and resource-intensive for training and deployment (Dhar et al., 2021), posing significant challenges for data collection and edge deployment in human sensing (Sobin, 2020; Liu et al., 2019).
Unlike other data modalities (e.g., vision or text), collecting human sensing data for training can be costly and even impractical. This is uniquely due to the data that presents brittleness (i.e., noise sensitivity, low SNR), inconsistency (i.e., user variance, scenario changes), and heterogeneity (i.e., differences in hardware configuration) (Lane & Georgiev, 2015; Jeyakumar et al., 2019; Teh et al., 2020). Collecting such data typically requires ethical approvals, as it involves sensitive information (e.g., location, motion patterns) that raises privacy concerns (Zhang et al., 2022). Manual annotation of large volumes of sensor data is labor-intensive and incurs high costs (Song et al., 2023). Importantly, human sensing spans various modalities (e.g., IMU, ultrasound, mmWave radar) and data types (e.g., spectrograms, Doppler profiles, time series), further amplifying the costs of applying DL solutions and naturally leading to cross-modality scenarios (Xiao et al., 2022; Zhang et al., 2023a).
Prior research has explored Few-shot Learning (FSL) (Wang et al., 2024), transfer learning (Dhekane & Ploetz, 2024), and their combination (i.e., cross-domain FSL) (Thukral et al., 2025; Yin et al., 2024) to notably reduce data collection costs by adapting pre-trained models to sensing applications. However, they still rely on large-scale labeled datasets from the same modality to costly train source models from scratch (Song et al., 2023), or face challenges such as modality shift and poor adaptability when leveraging pre-trained models (Oh et al., 2022; Tao et al., 2022; Guo et al., 2022), particularly in cross-modality settings. Recent advances in multi-modal learning (Yang et al., 2023; Chen & Yang, 2025; mengxi Chen et al., 2024; Madaan et al., 2024; Yang et al., 2024; Baldenweg et al., 2024) (e.g., CLIP-based models (Girdhar et al., 2023; Wu et al., 2023; Moon et al., 2023)) and cross-modal learning (e.g., image-to-sensor distillation (Zhao et al., 2018; Song et al., 2022; Gurbuz et al., 2020; Zhang et al., 2023c) or alignment-based transfer (Kamboj et al., 2025; Chen et al., 2024)), demonstrate the feasibility of knowledge transfer across modalities, but they remain modality-specific (i.e., relying on shared semantic feature spaces), and typically require large-scale paired data with resource-intensive training and deployment, leading to high costs and making them impractical for few-shot adaptation to support human sensing at the edge. These limitations motivate the need for a scalable and adaptable model transfer that can broadly adapt pre-trained models across modalities (e.g., vision or text) to new sensing modalities (e.g., radar or bio-signals) with few data.
In practice, however, addressing this need is uniquely challenging. Due to substantial differences in data characteristics—such as shape, distribution, and feature representation—between source and target modalities (i.e., modality shift), transferred models often suffer from significantly degraded performance (i.e., negative transfer) (Guo et al., 2022; Oh et al., 2022). Our experiments reveal that state-of-the-art (SOTA) baselines (Table 1) fail to reach oracle performance (Gong et al., 2019) and suffer from substantial overfitting when applied to human sensing tasks with few data, falling notably short of the goal, as shown in Figure 2(a). These highlight the underlying challenge of latent feature misalignment across modalities. Also, these methods often neglect resource constraints, rendering them impractical for edge deployment, especially in advanced multi-source settings (Zhao et al., 2020; Yue et al., 2021a; Lee et al., 2019). While methods such as neural architecture search (NAS) (Wen et al., 2023; Han et al., 2021), pruning (Frankle & Carbin, 2019; Shen et al., 2022), or quantization (Polino et al., 2018), aim to reduce resource demands, they often fail to maintain performance under modality shift and few data. As shown in Figure 2(c), our experiments using pruning (Shen et al., 2022) demonstrate significant model accuracy loss by 50.5% on average.
Figure 1: Overview. XTransfer transfers source models across modalities with few sensor data through model repairing (SRR pipeline) and layer recombining (LWS control). LWS control first segments source models into layers and operates layer-wise search across pools. At each pool, the pre-search check decides which layers need repairing, then SRR pipeline repairs them and LWS control selects layers of interest. These layers are incrementally recombined during the search, resulting in a compact model for enabling human sensing at the edge. Subfigures (a)–(c) illustrate the feature space evolution before and after repairing.
Driven by the increasing availability of high-quality pre-trained models from public repositories (e.g., PyTorch Hub (Hub, 2022)), we propose XTransfer, a first-of-its-kind method enabling modality-agnostic, few-shot model transfer. It flexibly leverages public pre-trained models and transfers knowledge across different modalities (e.g., image, text, audio, or sensing) to human sensing tasks using only few sensor data, significantly reducing the costs of large-scale data collection and training from scratch. Its resource-efficient design further optimizes edge deployment, enabling both scalability and adaptability to diverse human sensing applications. XTransfer uniquely repairs, searches and recombines layers of interest from pre-trained models under cross-modality FSL settings (see Appendix A), resulting in compact models. To mitigate modality shift and prevent overfitting, it enables model repairing using a novel Splice-Repair-Removal (SRR) pipeline (see Section 4). SRR pipeline transforms a few sensor data to broadly adapt pre-trained layers by repairing layer-wise misalignment and harmonizing latent feature distributions across modalities. XTransfer also features layer recombining using a novel Layer-Wise Search (LWS) control mechanism (see Section 5) to identify and recombine layers of interest by conducting a layer-wise search. Building on the insight that not all repaired layers contribute, LWS control discards such layers and recombines only useful ones to optimize model compactness using principles from NAS (White et al., 2023). To further accelerate the search in multi-source settings, a pre-search check strategy is designed to efficiently avoid unnecessary repairing.
The main contributions are summarized as: 1) a first-of-its-kind method, XTransfer, that enables modality-agnostic, few-shot model transfer with resource-efficient design for human sensing on edge systems, as shown in Figure 1; 2) a novel SRR pipeline that effectively mitigates modality shift and prevents overfitting, harmonizes latent feature distributions across modalities; 3) a novel LWS control mechanism that efficiently searches layers of interest from multiple sources in a layer-wise manner to create compact models; and 4) extensive experimental evaluations (see Section 6) under cross-modality FSL settings, benchmarking SOTA baselines (Table 1) on 8 public datasets with corresponding pre-trained source models across diverse modalities (e.g., image, text, audio, and sensing) and 7 target datasets (Table 2) using 3 commercial edge devices. The results show that XTransfer achieves a significantly enhanced accuracy-to-resource (ATR) ratio (see Section 6.1), outperforming single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively, while reducing layer search time by 2.1 to 4 times and substantially lowering the costs of data collection, model training, and edge deployment, demonstrating broad adaptability.
Figure 2: Preliminary study. (a) reveal baseline performance gap  [^2] We test a pre-trained ResNet18 on miniImageNet as source and HHAR dataset as target under a 5-shot setting (see Appendix A).. (b) shows the average similarity and FSL difficulty across all sensing datasets to source modalities (e.g., Image, Text, Sensing) using default reshaping (see Appendix B.1). 2 distinct areas represent similarity levels (A–hard, B–normal). Key findings: 1) compared to CUB, similarity levels across modalities are notably low, e.g., Text and Sensing fall into Area A, indicating a significant modality shift; 2) compared to Image-Unpair (i.e., no class pairing) in Area A, Image surprisingly falls into Area B, indicating that pairing classes may enhance cross-modality similarity (motivating our techniques in Section 4.1); 3) compared to Text and Sensing, Image exhibits more stable standard deviations and lower FSL difficulty, suggesting better potential for broad adaptability. (c) shows model pruning bottlenecks.
2 Related Work
Learning with few data for human sensing. Recent works in FSL and cross-domain FSL have proposed single- and multi-source methods. Single-source methods align latent features between source models and target data via distance-based objectives (Zhao et al., 2021; Yue et al., 2021b), whereas multi-source methods leverage the Wisdom of the Crowd principle through unsupervised source distillation (Zhao et al., 2020; Yue et al., 2021a) or maximal correlation analysis (Lee et al., 2019). Recent efforts have also explored fine-tuning LLMs (e.g., GPT2 (Zhou et al., 2023)) for time-series analysis. However, testing on benchmarks (Liang et al., 2021; Oh et al., 2022), our results reveal that they suffer from severe overfitting under cross-modality FSL settings. LLM-based data augmentation (Leng et al., 2024; 2023) helps reduce data collection costs, but remains modality-specific and orthogonal to our focus. Existing attempts on human sensing tasks (Gong et al., 2019; Wang et al., 2024; Thukral et al., 2025; Yin et al., 2024) remain constrained by a strong reliance on learning in the same modality, resulting in costly source data collection and training with limited adaptability. In contrast, XTransfer enables layer-wise model repairing to safely mitigate modality shifts using only few sensor data, achieving significant cost reduction with broad adaptability.
Resource constraints optimization. Model compression techniques such as model pruning (Frankle & Carbin, 2019; Shen et al., 2022), quantization using low-precision data types (Polino et al., 2018), model merging by fine-tuning shared layers (Padmanabhan et al., 2023), one-size-fits-all model (Cai et al., 2020), and NAS techniques such as model structure optimization (Wen et al., 2023; Han et al., 2021) or search using few data (Eustratiadis et al., 2024; Xu et al., 2022), have shown success in reducing model resource overhead. However, they strongly rely on either sufficient target datasets or minimal modality shifts to maintain model performance. XTransfer uniquely applies layer recombining to optimize model structures while strengthening layer-wise dependence, by efficiently searching and recombining layers of interest from pre-trained source models at scale. It not only creates compact models for streamlined edge deployment but also further enhances model repairing.
3 Preliminary
Figure 3: Design insights. (a) Layer-wise accuracy convergence using baselines is disrupted due to modality shift. (b) A notable MMC shift emerges and grows with increasing layer index, i.e., accuracy increases while MMC shift drops in Area A and begins to drop as MMC shift largely grows in Area B. (c) After repairing, layer S-score improves, but stagnation occurs at certain layers.
Insights in cross-modality FSL. We apply a standard cross-domain FSL benchmarking (Oh et al., 2022) to gain insights. It uses Earth Mover’s Distance (EMD) to measure similarity between source models from 3 modalities (i.e., Image, Text, Sensing) and targets including 5 human sensing and 2 well-known image datasets (Table 2). It also assesses FSL difficulty by fine-tuning source models on targets. Since source models involve a large number of classes  [^3] The number of classes in the source model should be greater than that in the target. (e.g., 100 classes in miniImageNet), we propose a class pairing strategy that pairs source and target classes exhibiting similar latent feature distributions across modalities. The key insights are shown in Figure 2(b).
Layer-wise analysis. We perform a layer-wise analysis to gain insights, focusing on accuracy changes and latent feature distributions at each layer. Since effective training relies on learning discriminative features (Islam et al., 2021), we benchmark latent feature distribution clustering and use Mean Magnitude of Channels (MMC) (Luo et al., 2022) of each layer as the default metric to quantify layer-wise performance, as it offers a lightweight measure for channel importance. We define a single layer or a dependent layer block (i.e., L-Blocks in Table 3)—where layers are not structurally separable—as the default segmentation unit (i.e., L-unit). Figure 3(a) shows that negative transfer (Zhang et al., 2023b) disrupts accuracy convergence across layers (i.e., layer-wise dependence), causing layer-wise misalignment and accuracy drops. These misalignments correlate with MMC shifts, where greater shifts correspond to larger accuracy drops shown in Figure 3(b), indicating inconsistent latent feature interpretation during fine-tuning. To track MMC shifts, we essentially employ the Silhouette score (S-score) (Shahapure & Nicholas, 2020), which captures inter- and intra-class distances and reflects the discriminability of layer-wise latent features.
Challenges. We aim to minimize layer-wise MMC shifts under cross-modality FSL settings, but several technical challenges arise. First, it is hindered by structural and semantic gaps between sources and target data, where classes and semantics often differ. Second, the pre-trained feature space at each layer is highly complex due to the multidimensional and noisy nature of MMC across source models from different modalities. This complexity increases with layer depth, as the strong layer-wise dependence formed during source model training becomes increasingly rigid, making it difficult to repair layer-wise misalignment using only few sensor data. Third, Figure 3(c) shows that S-scores may stagnate or even drop at certain layers (i.e., ineffective layers), impairing overall convergence. It indicates that not all repaired layers contribute positively, motivating our LWS control design to selectively recombine layers of interest while discarding others. The selection must also account for on-device resource constraints to control layer growth, while the rapidly expanding search space with increasing layers and source models further challenges the layer search.
4 Model Repairing
Enabled by SRR pipeline, this process safely repairs layer-wise misalignment by minimizing MMC shifts, effectively restoring layer performance from source models using few sensing inputs under cross-modality FSL settings. Three stages are designed, as shown in Figure 1. In the Splice stage, default reshaping (see Appendix B.1)—using fixed down/up-sampling—cannot be fine-tuned and may lose important features, especially when layer shapes vary during layer-wise search. We propose a compact connector placed between sensing inputs and layers, consisting of a Pre-header (adaptive convolutional layer), Resizer, and one encoder-decoder pair to bridge cross-modality structural gaps. In the Repair stage, each connector is fine-tuned by the generative transfer module to generate adaptive channels, preserve key sensing features, and minimize MMC shifts. Finally, the Removal stage applies channel removal (see Appendix B.2) to further enhance layer-wise performance.
4.1 Repairing in complex feature space
Since the original latent features from source models are well-trained for high discriminability, our intuition is to safely fine-tune few sensor data by aligning the misinterpreted latent features with the original features (i.e., anchors) to minimize MMC shifts. We hence propose cross-modality anchor-based alignment, which aligns source MMCs as anchors (i.e., anchor MMCs) with sensing MMCs. To handle complex feature spaces, our insight is to reduce MMC dimensionality by preserving key channels and excluding irrelevant or noisy ones.
Reduced-orthogonal feature space. We project anchor MMCs into a lower-dimensional, orthogonal feature space using principal component analysis (PCA) (Valpola, 2015) (i.e., anchor PCA space). Unlike the raw MMC feature space, where channel dimensions are often noisy and correlated, PCA enforces orthogonality across components, removing redundancy and suppressing noise. This enables a compact subspace (i.e., defaulting to 2 components) where the projected data captures maximum variance and forms highly clustered distributions. The optimized component weights (i.e., projection coefficients) highlight the most important channels contributing to layer performance. Given high discriminability of the original layers, the projected anchor MMCs preserve high clustering performance, with class distributions tightly centered (i.e., Anchor-center), as shown in Figure 1(a). Once initialized for each layer, this anchor PCA space is reused to project sensing MMCs.
Feature space alignment. After projecting sensing MMCs, we observe that the sample distribution for each class, along with their centroid (i.e., Sensing-center), appears disorganized compared to the well-clustered Anchor-center at each layer, as shown in Figure 1(a). Specifically, the scale and orientation of Sensing-center are notably different from those of Anchor-center, caused by MMC shifts. To align the scale, we first define a scale function based on the mean shift in inter-class distance $\mathrm{InterD}(\cdot)$, formulated as $S={{Mean}({InterD}(Pro(fs_{ij})))}\mathbin{/}{{Mean}({InterD}(Pro(ft_{ij})))}$, where $Pro(\cdot)={PCA}({MMC}(\cdot))$ denotes the space project function, $fs$ and $ft$ represent the latent features from both source and sensing modalities, respectively. $i$ denotes the index of pre-trained models, and $j$ denotes the layer index in each model $i$. Next, to align the orientation, we design a rotation alignment function that minimizes the cosine angle (Nguyen & Bai, 2010) by fine-tuning a multi-dimensional rotation matrix $M_{rot}]$ with dimensions matching the number of PCA components:
$\begin{split}\arg\min_{[M_{rot}]}Cos(Cen(Pro(fs_{ij})),Cen(Pro(ft_{ij})*S)\cdot[M_{rot}])\end{split}$ (1)
where $Cos(\cdot)$ denotes the cosine similarity function, and $Cen(\cdot)$ calculates the centroid of the projected distribution. Figure 1(b) shows the aligned results.
Anchor class pairing. Building on the key finding that class pairing boosts latent feature alignment across modalities (see Section 3), we propose selecting source classes with the highest S-scores as anchor classes to pair with sensing classes to facilitate the anchor-based alignment. We minimize pairing shift (i.e., the sum distance across each pair) via a linear sum assignment solved using the standard Hungarian algorithm, outputting the optimal pairing set $pairST$.
4.2 Repairing layer-wise dependence
Generative transfer module. To effectively restore layer-wise dependence, we design a generative transfer module that enhances the anchor-based alignment through generative learning. It consists of a generator (i.e., connector) and a discriminator—each frozen pre-trained layer paired with its anchor PCA space (Figure 1). Freezing preserves stable anchors and prevents overfitting. For each layer, given the optimal pairing set $pairST$, the generator is fine-tuned using the discriminator’s loss to minimize MMC shifts, and uniquely transforms sensor inputs to align with anchors.
Anchor-based repair loss. Building on the generative transfer module, we propose an anchor-based repair loss function ($loss_{srr}$) to harmonize latent feature distributions across modalities. Given each layer’s anchor PCA space, our objective is twofold: 1) minimize the distance ($D(\cdot)$) between the centroid of projected anchor MMC distribution and that of the projected sensing MMC distribution for each class pair $(c_{s},c_{t})\in pairST$ (i.e., positive loss), and 2) maximize the distance between sensing MMC samples with different labels $(c_{t}^{-})$ (i.e., negative loss). Since the projected anchor MMCs are highly clustered, they can provide the maximum margin ($M_{max}$) between anchor clusters. To further improve the negative loss, we update the function by calculating the anchor-based margin $margin^{c}=InterD(Pro(fs^{c}_{ij}))-IntraD(Pro(fs^{c}_{ij}))$, where the $M_{max}$ is selected, and $IntraD(\cdot)$ measures the intra-class distance. Hence, our loss function is formulated as:
$\begin{split}loss_{srr}=&\frac{1}{N}\sum^{N}_{(c_{s},c_{t})\in pairST}D(Cen(Pro(fs^{c_{s}}_{ij})),Cen(Pro(ft^{c_{t}}_{ij})))\\ +&\frac{1}{K}\sum^{K}_{(c_{t},c_{tk}^{-})}ReLU(M_{max}-D(Pro(ft^{c_{t}}_{ij}),Pro(ft^{c_{tk}^{-}}_{ij})))\end{split}$ (2)
where $K$ denotes the number of negative samples, and $N$ denotes the number of classes. The $ReLU(\cdot)$ is used to trigger the negative loss in the condition where the distance between the samples is less than $M_{max}$. Figure 1(c) shows the effectiveness of model repairing.
5 Layer Recombining
Working in conjunction with SRR pipeline for repairing each layer, XTransfer leverages layer recombining enabled by LWS control to: 1) perform the pre-search check that accelerates the search process, and 2) effectively select and recombine layers of interest to form a compact model.
5.1 Search control design
Following NAS principles (White et al., 2023), poor design choices (e.g., search space, actions, strategy) can disrupt the balance between model performance and resource efficiency. As layers of interest grow, resource overhead can increase exponentially, making dynamic search control essential under on-device resource constraints. To effectively manage, we define the search space as all layers from source models, structured into search pools containing layer candidates (i.e., repaired layers), each with width $I$ (number of models) and depth $J$ (number of L-units), forming a window of size $I\times J$. Search actions (init, continue, skip, cross) are defined to guide layer recombination within each designated search pool. Init selects the starting layer with a default fixed depth of 1. Continue and skip guide intra-model layer recombination, while cross enables inter-model layer recombination.
Search strategy under resource constraints. To effectively control search actions within each search pool, we propose a layer value function based on the S-score $Score(\cdot)$ to evaluate each repaired layer performance and the convergence between adjacent layers after each SRR process $SRR(\cdot)$. It is formulated as $V_{ij}=Score(SRR(Pro(ft_{ij}),MT_{ij},L_{ij}),Y_{t}))$, where $MT_{ij}$ denotes the connector and $L_{ij}$ represents each repaired layer. Under resource constraints, only the highest-valued repaired layer in each pool is selected. LWS control then checks whether its S-score exceeds that of the previous layer to maintain layer-wise dependence. If successful, the layer $L_{ij}$ is identified as the layer of interest and recombined, and the next pool window begins at layer index $j+1$. Otherwise, the current pool is discarded, and the search window shifts forward. Considering the constraints, we design a resource-constrained function that works in conjunction with $V_{ij}(\cdot)$, which is formulated as $RC(n)=exp(n/(L-2)-2)+1,n\in[1,~L]$, where $n$ denotes the current recombined layer index, and $L$ denotes the maximum number of L-units. Each layer’s adjusted overhead, $R(n)_{ij}$, is computed as $R(n)_{ij}=(Res(L_{ij})*RC(n))/Max(Res(P))$, where $Res(\cdot)$ calculates actual cost (e.g., FLOPs), and $Max(Res(P))$ is the peak cost in current pool $P$. The resource-constrained layer value function becomes $VR_{ij}=V_{ij}\mathbin{/}R(n)_{ij}$. The objective is defined as $\arg\max_{L_{ij}\in P}VR_{ij},\,s.t.\sum_{0}^{n}Res(L_{ij})\leq R_{Max}$, where $R_{Max}$ is device-specific resource budget.
5.2 Efficient search
Pre-search check. LWS control is designed to handle large search spaces from deep or multi-source models. While fine-tuning the connectors in SRR pipeline is efficient, the large search space can considerably slow down LWS control. We hence propose the pre-search check strategy to avoid unnecessary repair steps before applying SRR. We track S-score progression across three states—after anchor pairing (Anchor) (i.e., repairing objective), before SRR (Before), and after SRR (After) (i.e., repaired). SRR pipeline consistently improves S-scores from Before to After, approaching the Anchor level, indicating a strong correlation. Based on this, our intuition is to estimate repaired S-score (i.e., repair rate) of each layer based on Before and Anchor states to decide whether to bypass unnecessary repairs. Experiments show the actual repair rate grows exponentially with the recombined layer index (see Appendix B.4). We hence design a repair rate growth model formulated as $rate_{n}^{est}=exp(a*n)+b,n\in[1,L]$, where $a$ and $b$ are optimized via non-linear regression to minimize MSE to fit $rate_{n}^{est}$, with the objective function defined as $\arg\min_{a,b}\frac{1}{n^{{}^{\prime}}}\sum_{n=0}^{n^{{}^{\prime}}}(rate_{n}^{est}-rate_{n})^{2}$, where $n^{{}^{\prime}}$ is the number of recombined layers. The estimation begins after init. Once a new layer is selected, LWS control updates the estimation. Due to the limited updates, the estimation may be inaccurate. To mitigate, we also develop a dynamic search range mechanism that dynamically adjusts the search range in each pool (details in Appendix B.5).
6 Experiment
Table 1: State-of-the-art (SOTA) baselines.
  Baselines Group Description
  TL (Jiang et al., 2022) Single-source A standard transfer learning method by freezing all layers
FT (Oh et al., 2022) Single-source A standard fine-tuning method by updating all layers
ProtoNet (Chen et al., 2019) Single-source A distance-based FSL method
MAML (Chen et al., 2019) Single-source A learning-based FSL method
DAPN (Zhao et al., 2021) Single-source A cross-domain FSL method using adversarial ProtoNet
SemiCMT (Chen et al., 2024) Single-source A cross-modal method using contrastive alignment
GPT2 (Zhou et al., 2023) Single-(LLM) A large language model based method using fine-tuning
MDDA (Zhao et al., 2020) Multi-source A cross-domain FSL method using knowledge distillation
MCW (Lee et al., 2019) Multi-source A cross-domain FSL method using maximal correlation analysis
MetaSense (Gong et al., 2019) Oracle-(GT) An advanced FSL method based on MAML with sufficient target data
Table 2: Source and target datasets specifications.
  Dataset Modality Subject Class Type Shape Privacy
  Source
miniImageNet (Vinyals et al., 2016) Image - 100 - 3x84x84 Public
miniDomainNet (Zhou et al., 2021) Image - 126 - 3x224x224 Public
Office-31 (Saenko et al., 2010) Image - 31 - 3x224x224 Public
Office-Home (Venkateswara et al., 2017) Image - 65 - 3x224x224 Public
Caltech-101 (Fei-Fei et al., 2004) Image - 101 - 3x224x224 Public
Newsgroup (Thulasidasan et al., 2021) Text - 20 - 100x1x1000 Public
VoxCeleb (Nagrani et al., 2017) Audio - 100 Time-S 1x512x300 Public
OPPORTUNITY (Roggen et al., 2010) IMU 12 11 Time-S 77x1x100 Public
Target
HHAR (Stisen et al., 2015) IMU 9 6 Time-S 6x1x256 Public
WESAD (Schmidt et al., 2018) ECG 15 3 Time-S 10x1x256 Public
Gesture IMU/PPG 10 8 Time-S 7x1x200 Private
Writing Ultrasound 10 10 Time-S 1x1x256 Private
Emotion mmWave 10 7 Spectrum 3x224x224 Private
BP mmWave 10 - Doppler 1x1x499 Private
ChestX (Wang et al., 2017) Image - 5 - 1x1024x1024 Public
  Backbone Layers L-Blocks GFLOPs Params Kernel Shape
  ResNet18 (Chen et al., 2019) 18 9 3.67 11.18M 2D 3x224x224
ResNet18-1D (Chen et al., 2019) 18 9 0.39 3.9M 1D 100x1x1000
Multi-Conv4 (Zhao et al., 2020) 4x5 4 0.3 1.15M 2D 3x32x32
Multi-ResNet10 (Zhao et al., 2020) 10x5 5 0.7 24.5M 2D 3x84x84
Multi-ResNet18 (Zhao et al., 2020) 18x5 9 18.35 55.9M 2D 3x224x224
Conv4 (Chen et al., 2019) 4 4 0.02 0.07 1D 1x1x500
Conv5 (Gong et al., 2019) 5 5 0.19 30.42M 1D By Input
GPT2 (Zhou et al., 2023) 4x6 6 0.41 82.9M 1D By Input
Table 3: DL model backbone specifications.
6.1 Experiment setup
We benchmark XTransfer against SOTA baselines, grouped into single- and multi-source methods (Table 1). Evaluation is conducted on 8 widely-used public datasets under cross-modality FSL settings (Table 2), using their corresponding public pre-trained models as sources, from diverse modalities (e.g., image, text, audio, and sensing) with various backbones such as ResNet18 from PyTorch Hub (Hub, 2022) (Table 3). To evaluate XTransfer, we develop 4 real-world applications and build testbeds on 3 commercial edge devices (e.g., smartwatch, smartphone, Raspberry Pi). Based on these, we collect 4 human sensing datasets (marked as Private in Table 2) as targets, from 40 participants in real-world settings, with ethics approval from the Human Research Ethics Committee of our institute. Additionally, we include 2 public sensing datasets and 1 image dataset as targets for standard benchmarking. We apply standard Leave-One-Out Cross-Validation (LOOCV) (Gong et al., 2019) for evaluating human sensing datasets, and adopt accuracy-to-resource (ATR) ratio, defined as $ATR=Accuracy\mathbin{/}(\alpha Norm(FLOPs))+(1-\alpha)Norm(Params))$ for evaluating resource-accuracy efficiency, where $\alpha$ is set to 0.5 by default (see Appendix A for details).
6.2 XTransfer performance
Figure 4: Ablation study evaluating the performance of model repairing and layer recombining.
Repair performance. We first evaluate how our Repair loss reduces MMC shift to enhance the S-score of each layer with different MMC dimensionalities in the SRR pipeline. We use HHAR and the standard pre-trained ResNet18 on miniImageNet in a 5-shot setting. The channel removal is disabled in this experiment. We test with 4 different levels of MMC dimensionality (i.e., number of channels) ranging from 64 to 512. We employ N-Pair loss (Zhang et al., 2023c) and Triplet loss (Weller et al., 2022) in FSL sensing tasks as baselines. We run 15 rounds for each setting by default throughout the evaluation. Figure 4(a) presents that our Repair loss consistently achieves the highest S-score of 0.2 on average, outperforming both N-Pair loss (0.05) and Triplet loss (0.07), across all levels of MMC dimensionality. It indicates that Repair loss effectively enhances layer performance.
Efficient search performance. We now evaluate LWS control performance regarding model accuracy and layer search efficiency. We employ HHAR as target, ResNet18 on miniImageNet and Multi-ResNet18 on 5 image datasets as sources shown in Table 2, with a default search depth of 3 in 3-, 5- and 10-shot settings. We also utilize MetaSense in Table 1 as the oracle baseline. We compare different configurations: ResNet18 (Single), Multi-ResNet18 (Multi) without efficient search enabled, and Multi-ResNet18 (Multi-Efficient) with efficient search enabled. As expected, Figure 4(b) shows that Multi-Efficient achieves a substantial time reduction ranging from 2.1 to 4 times compared to Multi in 5- and 10-shot settings, while outperforming the oracle baseline. Additionally, Multi-Efficient improves average accuracy by 7.8% over Multi across the three settings. While Multi-Efficient takes a longer search time than Single, it successfully achieves superior accuracy, highlighting the benefits of using multiple source models in XTransfer. The results also reveal that both Multi-Efficient and Multi achieve lower standard deviation on average compared to Single, suggesting that using multiple source models offers commendable search stability. In short, the results prove that the proposed efficient search significantly accelerates the layer search process and achieves better model accuracy.
Impact of different sources. Since the selection of pre-trained source models plays a crucial role in XTransfer, we evaluate how different source models from various modalities affect the accuracy and resource overhead of the output models. We use HHAR, 2 pre-trained ResNet18s on both miniImageNet (Image) and VoxCeleb (Audio), and 2 pre-trained ResNet18-1Ds on both Newsgroup (Text) and OPPORTUNITY (Sensing), as shown in Table 2. We keep the other experimental setups unchanged. Figure 4(c) shows that the output models using Image source achieve the highest accuracy of 70.5% on average, compared to that of 69.6%, 64.3%, and 67% achieved by the models using Text, Audio, Sensing as sources, respectively. This suggests that the latent features learned from Image have better discriminability than those from other modalities. Since XTransfer prioritizes accuracy, we select Image source models by default. More ablation study results refer to Appendix C.
6.3 Comprehensive comparison
Cross-modality FSL performance. We evaluate the overall performance of XTransfer regarding both accuracy and ATR across 5 sensing datasets and 1 image dataset as targets (Table 2). We compare XTransfer against a range of SOTA baselines shown in Table 1. We also mark our methods as Our-Single and Our-Multi, using pre-trained ResNet18 on miniImageNet and Multi-ResNet18 with 5 corresponding image datasets as sources. In addition, we evaluate the fine-tuning performance of GPT2-small for comparison and align the parameters with the work (Zhou et al., 2023). All methods on target datasets are evaluated under the same cross-modality FSL settings (see Appendix A), with MetaSense (Gong et al., 2019) serving as the oracle baseline, trained on Conv5 with sufficient target sensor data (Table 3). For consistency, we also train the oracle using ResNet18 on the Emotion dataset to match 2D kernels. As shown in Table 4 and 5, both our methods outperform the baselines regarding model accuracy and ATR, and successfully reach or surpass the oracle in model accuracy across all target datasets in 5- and 10-shot settings. Specifically, both Our-Single and Our-Multi achieve a significantly higher ATR on average, surpassing the existing single- and multi-source baselines by 1.6 to 29 times and 16.6 to 98 times, respectively. In 3-shot setting, our method achieves slightly lower accuracy than the oracle on HHAR, Gesture, and Emotion. This indicates that extremely low samples may not fully represent the test data distribution, leading to less stable feature alignment. Since the layers are reused from pre-trained sources, the performance also depends on their quality and semantic overlap with targets. Notably, the flexibility of XTransfer allows further gains in such low-shot cases by using additional source models with higher quality or closer relevance. In summary, the results show that XTransfer achieves SOTA accuracy while substantially reducing data and resource costs, providing a scalable and adaptable method for human sensing at the edge.
Training resource usage. We also examine the training resource usage of XTransfer by measuring GPU utilization (Util) in the cloud. We focus on the above baselines that require model re-training or adaptation for transfer, e.g., MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and SemiCMT (Chen et al., 2024)), excluding methods that rely solely on fine-tuning or distance-based techniques. The experimental setup remains the same as above. As shown in Table 6, both our methods achieve the lowest GPU Util across all n-shot settings compared to baselines. Notably, both Our-Single and Our-Multi achieve a substantial reduction in GPU Util by 2.6 to 3.6 times and 2.5 to 3.5 times on average, respectively, compared to baselines. This is due to our layer-wise method, which progressively repairs and recombines only a subset of layers rather than training the entire model at once. It also highlights the effectiveness of our compact connector design. Despite leveraging multiple source models, Our-Multi shows no significant increase (e.g., only 3.67%) in training resource usage, further highlighting its efficiency and scalability. Hence, XTransfer significantly reduces cloud resource usage, enabling scalable model transfer.
HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
MetaSense (Gong et al., 2019) 63.2 69.0 75.2 55.1 64.0 68.3 66.3 73.4 79.4 72.1 83.3 89.6 55.5 56.3 65.5 24.8 28.1 31.7
ProtoNet (Chen et al., 2019) 50.6 45.7 49.8 43.9 49.3 52.0 50.8 55.8 59.2 69.5 78.7 73.4 51.3 49.2 54.8 21.7 23.4 24.5
DAPN (Zhao et al., 2021) 48.7 51.2 54.7 56.9 61.2 63.1 45.8 49.4 50.5 75.1 78.6 79.6 48.0 50.0 58.7 24.0 24.4 25.5
MAML (Chen et al., 2019) 36.6 42.3 42.4 39.6 57.9 65.6 37.2 41.3 39.4 31.6 39.7 35.2 22.3 26.9 31.1 22.5 20.4 21.9
SemiCMT (Chen et al., 2024) 33.3 38.9 48.9 37.3 40.0 44.4 23.5 34.4 42.5 27.3 38.6 48.0 24.0 33.6 46.7 25.3 25.6 28.0
GPT2 (Zhou et al., 2023) 41.0 34.0 45.0 48.0 35.0 36.0 49.0 54.0 63.0 52.0 71.0 75.0 - - - - - -
Our-Single 59.0 71.8 80.7 77.9 78.4 81.8 45.2 69.6 77.8 72.8 87.0 91.3 43.6 55.6 61.4 27.7 28.6 30.4
MDDA (Zhao et al., 2020) 13.2 17.7 18.8 40.0 39.3 41.9 11.2 13.9 15.2 12.5 10.9 9.0 14.4 15.3 16.0 20.6 19.6 19.9
MCW (Lee et al., 2019) 27.6 25.4 26.9 30.7 32.7 32.4 22.3 23.4 25.6 32.3 31.6 30.2 25.3 27.2 26.5 22.4 21.5 22.8
Our-Multi 59.3 74.3 78.4 76.4 77.8 78.7 48.7 73.1 78.6 74.1 86.1 90.4 42.3 55.1 58.6 29.0 30.0 30.2
Table 4: Comparison in model accuracy. Top-2 results are highlighted in bold.
HHAR WESAD Gesture Writing Emotion ChestX
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
ProtoNet (Chen et al., 2019) 0.51 0.46 0.50 0.44 0.49 0.52 0.51 0.56 0.59 0.69 0.79 0.73 0.51 0.49 0.55 0.22 0.23 0.25
DAPN (Zhao et al., 2021) 0.49 0.51 0.55 0.57 0.61 0.63 0.46 0.49 0.51 0.75 0.79 0.80 0.48 0.50 0.59 0.24 0.24 0.26
MAML (Chen et al., 2019) 0.37 0.42 0.42 0.40 0.58 0.66 0.37 0.41 0.39 0.32 0.40 0.35 0.22 0.27 0.31 0.22 0.20 0.22
SemiCMT (Chen et al., 2024) 0.33 0.39 0.49 0.37 0.40 0.44 0.24 0.34 0.43 0.27 0.39 0.48 0.24 0.34 0.47 0.25 0.26 0.28
GPT2 (Zhou et al., 2023) 0.11 0.09 0.12 0.13 0.09 0.09 0.13 0.14 0.17 0.14 0.19 0.20 - - - - - -
Our-Single 1.42 1.57 2.09 2.76 2.61 2.51 1.55 1.90 1.76 2.11 1.88 1.84 0.82 1.07 1.26 0.83 1.04 1.24
MDDA (Zhao et al., 2020) 0.03 0.04 0.04 0.08 0.08 0.08 0.02 0.03 0.03 0.03 0.02 0.02 0.03 0.03 0.03 0.04 0.04 0.04
MCW (Lee et al., 2019) 0.06 0.05 0.05 0.06 0.07 0.06 0.04 0.05 0.05 0.06 0.06 0.06 0.05 0.05 0.05 0.04 0.04 0.05
Our-Multi 1.71 1.47 1.64 3.00 2.70 2.75 1.20 1.25 1.63 1.93 1.96 1.53 0.96 1.23 1.73 0.68 1.24 0.83
Table 5: Comparison in ATR ratio. Top-2 results are highlighted in bold
HHAR (%) WESAD (%) Gesture (%) Writing (%) Emotion (%) ChestX (%)
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
DAPN (Zhao et al., 2021) 48.85 49.19 51.79 44.62 49.49 48.81 42.73 48.71 50.59 43.88 56.63 57.51 44.67 59.73 63.12 46.66 47.32 48.48
MAML (Chen et al., 2019) 59.59 79.32 89.72 52.58 70.27 77.52 53.11 66.75 73.21 72.56 85.47 92.22 63.39 72.78 75.01 44.52 55.46 61.36
SemiCMT (Chen et al., 2024) 55.60 73.43 78.91 38.83 50.33 67.89 57.02 72.16 98.04 75.14 79.12 81.79 60.30 63.43 68.60 41.54 50.05 63.43
Our-Single 18.27 18.90 19.48 15.95 16.46 16.98 20.67 21.71 22.51 19.92 20.67 21.08 18.24 19.20 19.96 16.97 17.58 18.80
Our-Multi 18.31 20.77 22.06 18.00 19.42 23.10 20.49 22.49 24.49 18.68 19.25 21.69 16.45 18.18 19.85 15.33 17.69 19.62
Table 6: Comparison in GPU utilization of each method during training. Top-2 results are highlighted in bold
7 Conclusion
This paper presents XTransfer, a novel method that enables modality-agnostic few-shot model transfer with resource-efficient design for human sensing applications. XTransfer harnesses the power of pre-trained models as free sources and requires only few sensor data to create compact models. It significantly reduces the costs associated with sensor data collection, model training, and edge deployment, offering a scalable and adaptable method for advancing human sensing.
Appendix
The appendix is organized as follows:
  • $\bullet$
    Section A illustrates more details on our experiment setup, including implementation details, developed real-world sensing applications, experimental setups, dataset statistics, evaluation metrics, and computation resources we used for our experiments.
  • $\bullet$
    Section B introduces additional technique details of different components in XTransfer.
  • $\bullet$
    Section C gives more experimental results about ablation study and benchmarking.
Appendix A Experiment setup details
A.1 Settings for learning with few data
A typical FSL method consists of two stages–meta-training and meta-testing (Chen et al., 2019). In meta-training stage, FSL requires a large-scale source dataset from the same modality (i.e., typically specific to the application) to train source models from scratch. The source dataset is divided into a support set and a query set  [^4] The support set and query set are assumed to be drawn from the same data distribution (Triantafillou et al., 2020).. Data for each set is loaded according to the n-way (i.e., number of classes) and n-shot (i.e., number of samples per class) format (Song et al., 2023). The meta-testing stage follows the same structure but uses unseen target data. Specifically, the target support set is used to adapt the source models, which are then evaluated on the target query set. Leave-One-Out Cross-Validation (LOOCV) is widely used as the de facto standard for evaluating datasets involving multiple users in human sensing (Gong et al., 2019). LOOCV works by leaving one user out as the target for meta-testing while using the remaining users to train the source models during meta-training in each round. The MetaSense (Gong et al., 2019) is used as an oracle baseline under FSL settings, where sufficient target sensing datasets are provided during meta-training stage. Unlike FSL, cross-domain FSL focuses primarily on meta-testing and transfers pre-trained source models across domains (i.e., different users), while remaining within the same modality (Oh et al., 2022). In this work, the cross-modality FSL settings are defined similarly to cross-domain FSL, but extend it to transfer pre-trained source models across different modalities.
A.2 Setups in preliminary study
As shown in Figure 2(a), we test a standard overfitting metric, defined as $\frac{\left|Accuracy_{train}-Accuracy_{test}\right|}{Accuracy_{train}}$ (López et al., 2022). In Figure 2(b), we test 3 source models (using a pre-trained ResNet18 from PyTorch Hub (Hub, 2022)) from 3 modalities, e.g., miniImageNet (Image), Newsgroup (Text), and OPPORTUNITY (Sensing), shown in Table 2. We also use 5 sensing datasets (HHAR, WESAD, Writing, Emotion, and Gesture) as targets, and employ 2 well-known image datasets (CUB and ChestX) as reference targets to the Image, detailed in Table 2. We use a default 5-shot cross-modality FSL setting with LOOCV enabled.
A.3 Setups in layer-wise analysis
In Section 3, we conduct a layer-wise analysis to gain insights into cross-modality model transfer with few sensor data. To set up, we continue to use the standard ResNet18 (Hub, 2022) and segment it as 9 L-Blocks, as shown in Table 3. We also employ the default reshaping (see Section B.1) on HHAR as target with 5-shot setting, and use the original source model input (miniImageNet) as reference, marked as Original. We examine the single-source baselines, including Transfer Learning (TL) (Jiang et al., 2022), Fine-tuning (FT) (Oh et al., 2022), MAML (Chen et al., 2019), DAPN (Zhao et al., 2021), and MetaSense (Gong et al., 2019) as the oracle baseline, shown in Table 1. To precisely track the MMC shift changes at each layer during fine-tuning,
A.4 Layer-wise metric
To precisely track the MMC shift changes at each layer during fine-tuning, we essentially use the Silhouette score (S-score) (Shahapure & Nicholas, 2020) as the layer-wise metric to evaluate how well the features are being fine-tuned. The S-score is in the range from -1 to 1 and being better if closing to 1. To verify whether the S-score indicates the layer performance, we observe the changes in the S-score and accuracy at each layer using the TL (Jiang et al., 2022) and Original methods. Figure 6(a) also shows that S-score and accuracy by Original are positively correlated, converging in a layer-wise manner. TL’s S-score also performs equivalently to its accuracy (i.e., S-score drop along with accuracy drop) in detecting the negative transfer across different layers.
A.5 Implementation
We implement XTransfer using PyTorch version 1.4.0. We also benchmark XTransfer with a full development cycle of cross-modality FSL tasks, including pre- and post-processing, training, and inference on commercial edge devices. Once pre-trained source models and few sensor data are processed in the cloud via XTransfer, compact models can be seamlessly converted to various DL platforms (e.g., PyTorch Mobile, TFLite) for broader edge deployments. To enable reproducibility of the results, we fix all random seeds in the code, set the evaluation mode consistently when evaluating output models, and enable CUDA convolution determinism. To measure model resource overhead (e.g., FLOPs and parameters), we use the standard PyTorch OpCounter. For regression tasks, class-based transformation (Shi et al., 2022) enables compatibility.
A.6 Developed sensing applications
To evaluate XTransfer on real-world human sensing targets, we essentially develop 4 human sensing applications, as shown in Figure 5. These include: (1) emotion recognition via facial expressions (i.e., Emotion) and (2) non-contact blood pressure monitoring (i.e., BP) based on an embedded mmWave radar setup, (3) handwriting tracking (i.e., Writing) using ultrasound based on a smartphone, and (4) gesture recognition (i.e., Gesture) using IMU and PPG sensors based on a smartwatch. The data collection is approved with ethics approval and involves 40 subjects. The corresponding target sensing datasets are shown in Table 2 with detailed specifications.
A.7 Hardware setup and computation environment
We evaluate XTransfer on real-world testbeds using 3 commercial edge devices, ranging from smartwatch to Raspberry Pi (Table 7). For example, Figure 5(a) presents an embedded mmWave radar setup which consists of a TI IWR1843Boost and a Raspberry Pi 4 using ROS Melodic on Ubuntu 18.04.4. In addition, we implement both smartwatch and smartphone testbeds using PyTorch Android version 1.10.0. We also employ servers equipped with NVIDIA GeForce RTX 3090 GPU to build our cloud environment.
A.8 Training details
In the Repair stage, we fine-tune the generative transfer module (see Section 4) for 100 episodes using SGD with a learning rate of 1e-2 and momentum 0.95. We apply a step-based learning rate scheduler with a step size of 20 and a decay factor of 0.5. We also use standard early stopping with a patience of 20. We retain the training parameters of the baselines in Table 1 as specified in their original implementations.
Figure 5: (a) Embedded mmWave radar testbed setup; (b)-(e) Built human sensing applications across different real-world settings.
  Device ROM RAM CPU Battery OS
  Galaxy Watch 4 16GB 1.5GB Exynos W920 247mAh Wear OS 3.5
Raspberry Pi 4 64GB 8GB Broadcom BCM2711 - Ubuntu 18.04.4
Redmi K20 Pro 128GB 8GB Snapdragon 855 4000mAh Android 11
Table 7: Edge devices specification.
Appendix B Technical details
Figure 6: Design insights. (a) shows layer-wise metric correlation. (b)(c) present efficient search insights into LWS control using multiple source models.
B.1 Default reshaping
To align with source model input shape, we develop a default reshaping to transform sensor data shape. It uses bilinear interpolation (Mastyło, 2013) (i.e., Resizer) to resize the height and width, and a fixed convolutional layer (i.e., Fixed header) with a $1\times 1$ kernel to match the input channels (Szegedy et al., 2014), placed between source models and sensor data.
B.2 PCA-based channel removal
In the Removal stage, channel removal is designed to further enhance layer-wise performance while optimizing resource usage. This is achieved by removing unnecessary channels at each layer. Existing pruning techniques (Shen et al., 2022; Frankle & Carbin, 2019) typically rely on channel importance metrics, such as the L2 norm, calculated from MMC. However, due to MMC shift, these metrics may inaccurately assess channel importance, resulting in degraded layer performance. To address this, we propose using PCA-based channel importance, derived from the optimized component weights during the repair process, as a more reliable metric for channel removal. Our objective is to remove unnecessary channels $\widetilde{ch}$ while preserving the maximum S-score ($Score(\cdot)$), formulated as $\arg\max_{(\widetilde{ch}\,\in CH)}Score(PCA(MMC(ft^{r}_{ij}),\widetilde{ch}),Y_{t})$, where $CH$ denotes the set of all channels, $r$ denotes the repaired latent features $ft_{ij}$, and $Y_{t}$ denotes the target labels.
B.3 Setups for efficient search insights
To gain design insights of efficient search for LWS control (see Section 5), we observe a complete process where all layers are repaired using the HHAR dataset and a Multi-ResNet18 model (i.e., 5 source models with 45 L-unit), as shown in Table 3. Figure 6(b) shows the changes in the S-score across three states: after pairing anchors (i.e., Anchor), before and after applying the SRR pipeline (i.e., Before and After, respectively), in a layer-wise manner. The results show that the SRR pipeline successfully increases the S-score from Before state (i.e., lower bound) to align with the Anchor state (i.e., the objective). This indicates a strong correlation between the After state’s S-score and both Before and Anchor values. These insights lead to the observation for designing the repair rate growth model.
B.4 Observation for designing repair rate growth model
Our key idea is to estimate the repaired S-score (i.e., repair value) of each layer before initiating SRR processing, allowing us to decide whether to bypass unnecessary repairs. To empirically examine this, we first define a repair rate function to capture this correlation as $rate_{n}=(After-Before)/Anchor=(V_{ij}^{n}-Score(Pro(ft_{ij}^{n}),Y_{t}))/Score(Pro(fs_{ij}^{n}),Y_{s})$, where $Y_{s}$ denotes the source labels, and we normalize $rate_{n}\in[0,1]$. We then repeat the process multiple times to observe the changes in $rate_{n}$. Figure 6(c) shows that the repair rate demonstrates exponential growth. Thus, we design the repair rate growth model based on an exponential growth model (see Section 5.2).
B.5 Dynamic search range mechanism
Based on the estimated repair rate, LWS control can initially use a small, fixed search range to efficiently filter out most layer candidates before initiating SRR processing. However, because the updates to $rate_{n}^{est}$ are limited, the estimation may be inaccurate, potentially affecting selection decisions (e.g., missing the optimal layer candidate). To mitigate this risk, we develop a dynamic search range mechanism that adjusts the search range to balance accuracy and efficiency. This mechanism is formulated as $range_{n}=(1+\hat{u}^{\pm}rate_{n}^{est})*range_{n-1},\,range_{n}\in[0.2,1]$, where the range starts at 0.5 and adjusts based on the previous range. Here, $\hat{u}$ denotes a signed unit vector. If $rate_{n-1}^{est}$ is less than $rate_{n-1}$, it indicates that the estimation might be inaccurate and $\hat{u}$ will be positive to increase the range in searching for the optimal layer. Otherwise, the range will be reduced to improve search efficiency.
B.6 Post fine-tuning
After LWS control is completed, we add either a linear classifier or a regression layer (e.g., a fully connected layer) at the end of the model to finalize its backbone. To strengthen layer-wise dependence, we perform a quick post fine-tuning process, updating only the connectors and fully connected layers.
Appendix C Additional experiment results
C.1 Feature space alignment performance
In the process of feature space alignment (Section 4.1), we use a standard Hungarian algorithm based on default distance metrics to select and pair classes. To verify its effectiveness, we keep using the setup of HHAR, the standard pre-trained ResNet18 on miniImageNet, and the standard PCA under 5-shot settings to examine how our repair loss performs with/without the feature space alignment, especially in the initial layer repairing. Repairing the initial layer is critical, as it often introduces a large, disorganized feature space and significantly affects the repair of subsequent layers. In Table 8, the results show that with the alignment, the repair loss decreases considerably faster by 1.87 times on average, compared to the case without alignment. This indicates that the design of feature space alignment notably contributes to the repairing process.
Epoch (#) #1 #20 #40 #60 #80 #100
With Alignment 0.486 0.355 0.350 0.339 0.333 0.331
Without Alignment 0.476 0.406 0.393 0.383 0.378 0.387
Table 8: Feature space alignment performance across training epochs.
C.2 Impact of PCA variations
Since our method uses the repair learning mechanism that relies on backpropagation (BP) through the reduced-orthogonal feature space, standard PCA can project data into a linear subspace and maintain a closed-form, differentiable structure, which enables smooth integration with BP in our end-to-end learning pipeline. Compared to advanced choices such as kernel PCA, we find that it introduces a non-linear mapping into a high-dimensional feature space using a kernel function, and its projections are generally not differentiable with respect to input features. This makes it incompatible with our pipeline. To test whether linear PCA variations affect the overall performance of XTransfer, we compare standard PCA against sparse PCA using the setup of HHAR based on the ResNet18 under 3 few-shot settings with a default search depth of 3. As shown in Table 9 (#2 denotes 2-component setting), sparse PCA reduces learning time slightly by 11.6% on average, but suffers from a notable accuracy drop by 4.4% on average. These results support the choice of standard PCA as a balanced and practical solution.
PCA Variation 3-shot 5-shot 10-shot
Accuracy (%) Time (mins) Accuracy (%) Time (mins) Accuracy (%) Time (mins)
Standard PCA-#2 58.96 3.9 71.85 5.6 80.74 10.3
Sparse PCA-#2 58.52 3.8 64.67 5.3 75.11 8.8
Table 9: Impact of PCA variations on accuracy and training time.
C.3 Impact of PCA components
We then evaluate the impact of PCA components on the model accuracy and time efficiency of XTransfer. Since the feature space with different dimensionalities adjusted by PCA components may impact our method performance, we test 3 different levels of PCA components, including PCA-#2 (top two components), PCA-80% (selected components covering 80% of the feature variance), and PCA-Max (maximum components available). We continue to use the same setup as above with the standard PCA. The results in Table 10 show that as the number of PCA components increases, the training time rises notably. In fact, PCA-Max results in an average accuracy drop of 4.5% compared to PCA-80%. By contrast, PCA-#2 achieves the best accuracy-to-time ratio across all n-shot settings, with competitive accuracy and the lowest training time, hence we use PCA-#2 as the default throughout the evaluation.
PCA Components 3-shot 5-shot 10-shot
Accuracy (%) Time (mins) Accuracy (%) Time (mins) Accuracy (%) Time (mins)
PCA-#2 58.96 3.9 71.85 5.6 80.74 10.3
PCA-80% 63.61 5.2 74.72 7.6 80.00 12.9
PCA-Max 60.00 7.2 75.56 8.1 70.28 16.9
Table 10: Impact of PCA components on accuracy and training time.
Figure 7: Ablation study evaluating the performance of components, search parameters, and applications.
C.4 Removal performance
We next evaluate how channel removal enhances the average S-score of selected layers. We continue to use the same setup as above under 5-shot settings. We employ L2 norm-based (i.e., L2-based) model pruning (Frankle & Carbin, 2019; Shen et al., 2022) as the baseline and compare it with the setting where channel removal is disabled (i.e., Non-Pruning). Figure 7(a) demonstrates that the proposed channel removal outperforms both Non-Pruning and L2-based pruning across all settings, achieving an average S-score of 0.1 by selected layers. L2-based pruning achieves a lower average S-score than Non-Pruning, indicating that the channel importance may be inaccurately estimated, resulting in degraded performance. These results show that the proposed channel removal further improves layer performance by accurately removing unnecessary channels at each layer.
C.5 Impact of search depth
In this experiment, we investigate how different search depths affect the accuracy and resource overhead (e.g., FLOPs) of output models. Varying the search depth alters the search window sizes with the number of layer candidates, which can lead to different output model backbones and accuracy. To quantify these differences, we employ HHAR as target, ResNet18 on miniImageNet (Single) and Multi-ResNet18 (Multi) on 5 image datasets as sources (shown in Table 2) in 3-, 5- and 10-shot settings. We also utilize MetaSense in Table 1 as the oracle baseline. Both Single and Multi models have up to 9 L-unit, hence we set three different search depths as D1, D3, D9. Figure 7(b) shows that the output models of both Single-D1 and Multi-D1 result in high FLOPs, while that of both Single-D9 and Multi-D9 achieve low FLOPs. As search depth decreases, the search process becomes longer, and it triggers LWS control to make more decisions on recombining layers, which results in a suboptimal model backbone with high FLOPs. Although both Single-D3 and Multi-D3 do not produce the lowest FLOPs on average, they outperform others and achieve the highest average model accuracy of 70.52% and 70.69%, respectively, across the three settings. Therefore, we set D3 as the default search depth, as it provides the best balance between resource overhead and accuracy of the output models.
C.6 Impact of model backbones
The current implementation of XTransfer uses homogeneous pre-trained source models. Technically, the layer-wise repairing and recombination techniques impose no fundamental restrictions on the choice of model backbone and work with various architectural variants. To verify this, we evaluate the trade-off between accuracy and computational cost using different source backbones. We continue to use HHAR, and test Multi-Conv4, Multi-ResNet10, and Multi-ResNet18 in a 5-shot, multi-source setting shown in Table 3. The results in Table 11 demonstrate that the average accuracy declined, transitioning from Multi-Conv4 to Multi-ResNet10, and then increasing with Multi-ResNet18. Interestingly, Multi-Conv4, despite having fewer parameters, outperforms Multi-ResNet10, which may be attributed to differences in model quality or distributional divergence between source and target. Multi-ResNet18, though with higher FLOPs, achieves the best accuracy. It suggests that larger models trained on large-scale cross-modality datasets exhibit superior latent features, which notably contribute to such model transfer. Our techniques such as the connectors are designed to flexibly handle the layers with different shapes/structures, and can be extended to heterogeneous model backbones. In addition, our method focuses on latent feature alignment rather than specific feature extraction. It remains fully compatible with advanced source models (e.g., with temporal encoders), which can be seamlessly integrated into our pipelines with minimal or no modifications.
Backbone Accuracy (%) FLOPs (G)
Multi-Conv4 65.56 0.01
Multi-ResNet10 63.61 0.14
Multi-ResNet18 78.44 2.19
Table 11: Performance of different source model backbones in terms of accuracy and FLOPs.
C.7 Training efficiency
To test the practical training efficiency in the cloud, we further evaluate training time across the selected baselines. We use the same experimental setups as in Table 6, Section 6.3. For MAML (Chen et al., 2019) and SemiCMT (Chen et al., 2024), which require model re-training from pre-trained source models to adapt to the target modality classes, training time includes both re-training and fine-tuning. For DAPN (Zhao et al., 2021), it reflects model adaptation process on few target data. The results in Table 12 show that both Our-Single and Our-Multi achieve significant training time reductions compared to MAML. Our-Single also matches or exceeds DAPN and SemiCMT in training efficiency. Notably, even when using up to 5 source models, Our-Multi remains practical training efficiency, requiring only 2.05 times more training time on average compared to Our-Single across all datasets and shot settings. This further supports the practicability and scalability of our method.
HHAR (mins) WESAD (mins) Gesture (mins) Writing (mins) Emotion (mins) ChestX (mins)
Method 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot 3-shot 5-shot 10-shot
DAPN (Zhao et al., 2021) 2.38 4.62 11.78 1.49 2.54 6.78 3.36 6.12 15.72 4.12 7.71 20.97 2.99 7.34 18.86 2.14 4.57 12.04
MAML (Chen et al., 2019) 124.28 125.66 183.20 71.71 77.07 115.87 118.79 129.79 191.95 165.06 129.27 194.36 94.81 98.40 153.20 66.20 72.17 106.97
SemiCMT (Chen et al., 2024) 3.02 8.27 10.23 1.82 2.48 5.63 4.81 9.42 12.83 10.10 13.92 18.75 3.39 7.32 11.48 2.84 5.28 8.33
Our-Single 3.98 5.57 7.95 2.12 2.99 4.33 5.64 7.80 11.24 7.94 10.54 14.69 5.58 7.43 9.88 3.65 4.75 6.84
Our-Multi 9.25 15.33 15.25 4.40 5.38 7.21 12.49 15.12 24.37 15.35 17.36 32.38 11.72 13.39 19.75 8.41 9.33 13.61
Table 12: Comparison in training time of each method.
C.8 Regression task performance
We also evaluate the effectiveness of XTransfer in a regression task using the BP dataset, which includes both systolic (SBP) and diastolic (DBP) sensor data shown in Table 2 collected by the blood pressure monitoring application. We employ the pre-trained ResNet18 on miniImageNet as the source model and use the 5-shot setting. To integrate this into the SRR pipeline, we transform the regression task by treating individual regression labels as distinct classes for anchor class pairing. For benchmarking, we use mean error (ME) and standard deviation (STD), following the Advancement of Medical Instrumentation (AAMI) standard, which accepts an ME within 5 mmHg and STD within 8 mmHg (Stergiou et al., 2018). We employ a random forest (RF) based method (Shi et al., 2023) as the oracle baseline. Besides, we train a Conv4 backbone in Table 3 (Conv) using LOOCV, and the fine-tuned Conv (ConvFT) as the baselines. Figure 7(c) shows that both XTransfer and the RF methods surpass the AAMI standard. In particular, XTransfer outperforms the baselines and RF in estimating both SBP (ME: -0.64mmHg, STD: 7.2mmHg) and DBP (ME: 0.07mmHg, STD: 7.47mmHg) tasks. These results demonstrate that XTransfer is highly efficient in handling regression tasks with few sensor data.
C.9 On-device performance
[…]
Watch Pi Phone
Dataset Type FLOPs Params Latency (ms) Mem (MB) Latency (ms) Mem (MB) Latency (ms) Mem (MB)
HHAR Single 1.51G 0.75M 625.5 125.5 274.18 246.62 26.86 179.5
Multi 1.46G 2.16M 634.5 140 249.26 247.83 28.34 209
WESAD Single 1.19G 0.74M 604.5 120 240.05 244.64 25.77 168
Multi 0.89G 0.34M 460.5 122 209.52 235.55 21.57 181
Gesture Single 1.21G 2.25M 706.5 135 259.25 251.26 31.5 192
Multi 1.45G 2.09M 692 137 268.79 253.7 30.96 190
Writing Single 1.32G 0.87M 467.5 126 261.71 242.33 29.79 183.5
Multi 1.28G 0.73M 513 124.5 285.2 247.01 34.67 185
Emotion Single 1.51G 0.49M 767.5 130.5 390.36 271.02 41.1 171.5
Multi 1.43G 2.14M 781 142 285.16 251.48 37.9 189.5
ChestX Single 0.64G 0.72M 690.5 126 210.45 237.45 29.57 165
Multi 1.06G 0.7M 670 130.5 261.14 236.96 30.14 183.5
ResNet18 - 3.67G 11.18M 1060 189 703.52 303.75 89.8 311
Multi-ResNet18 - 14.68G 55.9M 5816 356 4242.27 482.26 454.44 436
Table 13: On-device Performance.